### v1

# VisLTR: Visualization-in-the-Loop Table Reasoning

Abstract
Table reasoning transforms user requirements into corresponding answers according to the provided table, which is often integrated with natural language interfaces for lay users to explore tabular data effortlessly. Recent research exploits large language models to facilitate table reasoning, by transforming vague user requirements into structured query languages (SQLs). However, these SQL-based approaches often overlook changes in data patterns, suffer from LLM drift, and limit exploration to only text queries. To this end, VisLTR is designed as a visualization-in-the-loop table reasoning framework that leverages visualizations as a proxy to provide concise data representations, capture interesting data patterns, and support cross-modal analysis. We describe VisLTR as a process consisting of four major modules: 1) visualization alignment that utilizes large vision-language models to align visualizations across various modalities, including chart, text, and sketch; 2) visualization referencing that decomposes a table into multifaceted visualization references that comprehensively represent the table; 3) visualization pruning that incorporates data and retrieval pruning to excise visualization references with poor information and enhance retrieval efficiency; and 4) visualization interaction that offers an interactive visual interface with multi-modal interactions for user-friendly table reasoning. Quantitative evaluation demonstrates the effectiveness of the alignment model in cross-modal visualization pairings. We further demonstrate applications of the framework on various table reasoning tasks such as table summarization and pattern detection.
keywords:
Table reasoning, visualization reference, multi-modal learning
[…]
VisLTR supports data change pattern identification and multi-modal explorations of text queries and hand-drawn sketches using visualizations as a proxy between a table and user requirements (bottom). The proposed approach advances existing SQL-based LLM methods for table reasoning that overlook changes in data patterns, suffer from LLM drift, and limit exploration to only text queries (top).
1 Introduction
Table reasoning is the task of deriving answers from tables (i.e., tabular data) based on specific questions[58], supporting a wide range of applications such as fact verification[6] and question answering[21]. The key challenge of table reasoning lies in understanding both the intent of queries (e.g., a text query) and the semantics of tables, and accurately generating answers that require logical reasoning. Existing methods for table reasoning can be broadly categorized into generic table reasoning that directly generate answers using pre-trained models or large language models (LLMs); and program-aided table reasoning that utilizes executable query languages (e.g., SQL or SPARQL) as the bridge between user queries and answers derived from the table[49]. Generic table reasoning performs reasoning implicitly[49, 5], which, however, is poor at tracking how answers are derived and increases the risk of plausible but incorrect responses. Instead, program-aided table reasoning ensures precise data retrieval and offers transparency in the reasoning process, which is essential for applications such as fact verification[56, 7]. Therefore, program-aided reasoning methods are attracting significant attention from both academia and industry[15].
Despite the growing popularity of program-aided reasoning methods, there exist several limitations. One of the primary issues stems from SQL’s inherent lack of analytical functionality[24], which leads to an overlook of data change patterns. Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A) illustrates a scenario where users seek to examine the price trend of Apple during March. Despite successfully retrieving a sub-table containing the price of Apple in March, SQL-based LLMs fall short when it comes to further querying and discerning change patterns within this sub-table. Second, LLMs are known to encounter drift issues, which refer to the loss of information in the middle of long contexts. LLM drift is common when processing large data tables[4], where the LLM’s reasoning process may encounter errors due to the loss of information in the middle, as illustrated in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(B). Last, table reasoning is often integrated with natural language interfaces, to facilitate lay users’ exploration of tabular data effortlessly. However, there are scenarios in which laymen are unable to express their intentions in natural language but would prefer to use hand-drawn sketches. An example is shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(C), where a user may not know the term “two-valley” but can easily sketch this pattern. Existing SQL-based LLMs do not support non-verbal interactions, which are highly demanded.
To overcome the limitations, we introduce a visualization-in-the-loop table reasoning framework that incorporates visualization references into large vision-language models (LVLMs) to complement SQL-based LLMs for table reasoning. Our key idea is to transform a tabular data into a set of meaningful and insightful visualizations, termed visualization references, which serve as a proxy to capture data insights and align with human intentions. This framework has several advantages compared to existing SQL-based table reasoning methods. First, visualization references can capture the short- and long-term data change patterns. Considering the price trend query as an example (Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A)), the retrieved visualization reference could capture the “two-peak” price trend of Apple during March. Second, regardless of the length of data facets, the encapsulated visualization references can capture and retain the data change patterns in fixed-size images. Adopting fixed-size visualization references as the representation of data facets can effectively circumvent the ‘drift’ issues encountered during reasoning. As Figure VisLTR: Visualization-in-the-Loop Table Reasoning(B) shows, the stock price change over a year can be represented as a single visualization, where the price bottom can be easily located in this reference. Moreover, leveraging visualizations as a proxy offers a new modality of hand-drawn sketches for interacting with the tabular data, which align more intuitively with human cognition[27]. When querying for a hand-drawn sketch as shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(C), it is easier to retrieve a similar visualization reference among the candidates than to retrieve a similar change pattern in the tabular data[35].
However, the strategy presents several challenges. First, data tables often contain a variety of long- and short-term patterns, which should be comprehensively captured by the visualization references. (C1) Ensuring the visualization references are comprehensive enough to accurately represent an input data table, thereby meeting the diverse analytical needs, poses the first challenge. Second, (C2) effectively aligning textual descriptions, visualization images, and hand-drawn sketches to bridge user queries with visualization references, meanwhile enabling multi-modal interactions presents a challenge. Existing models for vision-language tasks such as chart question answering (CQA)[19] are not specifically designed to meet this requirement, thus being less effective in identifying data change patterns. Last, to comprehensively represent a table, a large number of visualization references will be generated. (C3) Effectively indexing the visualization references and supporting efficient visualization retrieval for timely user response present a significant challenge.
To evaluate the effectiveness of the proposed framework, we design a prototype system, which encompasses four main modules aimed at addressing the challenges C1--C3. The visualization referencing module decomposes an input table into data facets and processes the facets into corresponding visualization references that comprehensively capture the essence of the table. To ensure a comprehensive representation of data patterns in tabular data, we adopt a multifaceted process to get meaningful facets and then transform each facet into a fixed-size visualization reference (addressing C1). Second, in the visualization alignment module, we fine-tune a large vision-language model (LVLM) to connect visualizations of various modalities including charts, text, and sketches (addressing C2). To fine-tune the model, we construct a new cross-modal visualization dataset using data augmentation for chart-text pairing and user labeling for chart-sketch pairing, taking into account user intentions. We further leverage the CLIP model[41] enriched with two-level cross-entropy loss and bidirectional hinge-based triplet ranking loss to enhance the alignment. Next, the visualization pruning module leverages visualization pruning to filter out less informative visualization references, and visualization indexing to enhance the efficiency of the retrieval process for the embeddings of visualization references stored in a vector database (addressing C3). Finally, we design a visualization interface that enables users’ exploration of time-series data tables through multi-modal interactions. Here, we incorporate a “decomposing-execution-filling” strategy to integrate the output generated by VisLTR with the input table. Specifically, a user query is initially decomposed into visualization retrieval tasks, which are executed in the vector database, and their corresponding text descriptions are extracted by the fine-tuned LVLM. Subsequently, these descriptions are filled back into the input table to generate textual responses for the user. Quantitative evaluations demonstrate the effectiveness of the fine-tuned LVLM, and case studies illustrate the system’s applicability in diverse application scenarios. Dataset, fine-tuned model, and evaluation results are released at https://osf.io/yun9d/.
In summary, we make the following contributions:
(1) Visualization-in-the-Loop Framework. We introduce a visualization-in-the-loop framework that leverages visualizations as the proxy to support table reasoning. The framework complements existing program-aided approaches for table reasoning, particularly in capturing data change patterns and facilitating non-verbal exploration.
(2) Multi-modal Visualization Alignment. We design an effective LVLM that learns a joint embedding space across three visualization modalities - charts, text, and hand-drawn sketches. The model aligns multi-modal user intentions with visualization references, allowing users to explore and analyze tables through various interactions.
(3) VisLTR System. We develop a prototype system that supports the exploration of large-scale and multiple time-series tables through intuitive user interactions. Various case studies and quantitative results prove the effectiveness and applicability of VisLTR.
2 Related Work
2.1 Table Reasoning
Table reasoning integrated with natural language (NL) interfaces supports lay users to explore tabular data effortlessly. This capability is essential for many complex tasks such as fact verification[6] and question answering[21]. Existing table reasoning methods can be categorized into generic table reasoning and program-aided table reasoning[49]. Generic reasoning methods employ LLMs to interpret and respond to free-form NL queries by generating answers directly from tables[5]. Due to the complexity of table structure, these methods face difficulties in dealing with large tables. Moreover, explainability is limited due to their end-to-end nature. To facilitate complex table reasoning, program-aided table reasoning utilizes executable query languages, such as SQL, to form a more structured and accountable bridge between user quires and the derived answers[11, 31, 20]. Recent research has mostly focused on leveraging LLMs to transform complex user utterances into SQL queries with data context awareness[56], using prompt engineering[25], chain-of-thought approaches[49, 57], and pre-training techniques[48]. However, SQL-based LLM table reasoning approaches encounter intricate limitations due to their reliance on SQL. First, SQLs primarily cater to numerical queries from tables but lack support of analytical functionality[24], overlooking data change patterns. Second, when confronted with large tables, LLMs may encounter drift issues[4] while accessing information in the middle of long contexts. Moreover, in many scenarios, relying solely on textual inputs restricts the expression of non-verbal intentions, such as hand-drawn sketches.
To overcome these limitations, this work introduces a novel framework that leverages visualizations as a proxy between tables and user queries, complementing SQL-based table reasoning by addressing challenges such as data change pattern identification and multi-modal interaction. The strategy is feasible, as visualizations can depict intuitive data patterns and illustrate complex underlying processes[14], facilitating swift and effective interpretation by the human brain[26]. However, this approach also presents challenges to be tackled. First, tabular data typically encompass numerous features, leading to a plethora of potential visualizations[54, 46] that complicate interactive interactions. Second, there are no direct approaches available to connect free-form user queries with visualizations.
Figure 1: An overview of the proposed framework, which mainly contains four main modules: 1) the visualization referencing module decomposes a raw data table into facets and maps them to visualization references; 2) the visualization pruning module leverages pruning and indexing techniques to ease the storage and improve the retrieve efficiency; 3) the visualization alignment module leverages a fine-tuned LVLM to enable alignment of various visualization modalities, including chart, text, and sketch; and 4) the visualization interaction module decomposes user queries into retrieval tasks with visualization references and utilizes the LVLM to generate the corresponding trend words to fill the textual responses for users.
2.2 Multi-modal Visualization Alignment
Given that visualizations are inherently stored as images, we opt to leverage emerging vision-language models to embed visualization images and facilitate multi-modal user interactions. In recent years, significant advancements have been made in utilizing vision-language models for chart-based tasks, such as CQA[36, 37, 19], chart retrieval[52] and generation[43, 53]. Though CQA models can provide answers for user utterances, these models only support text-chart interactions, lacking the capability for cross-modal interactions involving multiple modalities such as text, chart, and sketch preferred by laymen. Furthermore, these models are not specifically designed for effectively identifying data change patterns (see Sect. 6.1), a gap that this work seeks to fill in the context of table reasoning.
To this end, we choose to construct a novel vision-language model with an innovative visualization alignment strategy. Most existing multi-modal data alignment methods rely on contrastive language-image pre-training (CLIP)[41], which has demonstrated its adaptability and proficiency across various tasks[59, 29, 60], highlighting the effectiveness of joint embedding spaces. However, training an effective CLIP-based alignment model for visualizations is non-trivial. First, this task involves aligning more than two modalities, including text, hand-drawn sketches, and various types of charts, whereas the original CLIP model was designed to support text-image alignment. To encompass more modalities, we adopt joint embedding strategies that leverage charts as naturally aligned supervision and emergent alignments to unify other modalities[16]. This approach allows us to prepare paired data only for chart-text and chart-sketch, rather than other pairings such as sketch-text. Second, the effectiveness of CLIP is largely attributed to its extensive training on vast (natural) image-text datasets, whereas there is a lack of suitable training datasets for the task in this work. Efforts have been made to construct datasets for promoting deep learning models for chart-text (e.g.,[23, 36]), chart-chart (e.g.,[34, 28, 55]), chart-sketch (e.g.,[35, 45, 32]) similarity measurement. However, these different modalities are not inherently unified in one model, and many of these datasets are not specifically designed for identifying data change patterns. To address this gap, we utilize data augmentation and user labeling strategies to supplement existing datasets in training the multi-modal visualization alignment model.
2.3 Visualization Storage, Indexing, and Retrieval
To comprehensively represent tabular data, our approach will generate a large number of visualization references and encode them into vector representations, thus raising new challenges for efficiently storing, indexing, and retrieving them. With the recent advancements in Vector Database Management Systems (VDBMS)[39], there is potential to meet our demand. VDBMSs are specifically designed for managing high-dimensional data, such as embeddings, offering sophisticated storage, indexing, query processing, and query optimization capabilities[3]. Their ability to perform efficient similarity searches in high-dimensional spaces makes them particularly suited for tasks involving the retrieval of embeddings based on content or semantic similarity. This supports various applications such as information retrieval, e-commerce, and recommendation platforms[17, 47]. Existing VDBMS can be roughly categorized into two types: native systems and extended systems[39]. The former, such as Chroma[1] and EuclidesDB[2], are tailored for vector data management, while the latter, such as AnalyticDB-V[50], are built on top of an existing data management system.
In this work, we first introduce a visualization pruning method to filter out less informative visualizations, and effectively reduce storage demands while maintaining the quality of stored visualizations. We then integrate Chroma[1] into VisLTR and utilize an efficient graph-based index to store and index the large collection of visualization references. An Approximate k-Nearest Neighbors (ANN) algorithm is incorporated for quick and accurate retrieval of visualizations. The combination of pruning with indexing strategies allows for quick retrieval and handling timely user queries.
3 The Visualization-in-the-Loop Framework
The proposed visualization-in-the-loop framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Figure 1.
Visualization Alignment. As the cornerstone of our framework, an LVLM is fine-tuned to establish a single joint embedding space for various modalities, including chart ($\mathbb{C}$), text ($\mathbb{T}$), and sketch ($\mathbb{S}$). To address the issue of a lack of suitable training data, we first curate a new dataset with chart-text pairings augmented from existing Chart-to-Text dataset[23], and chart-sketch pairings through user labeling. We apply pre-trained CLIP[41] to fine-tune a multi-modal model on our composite dataset. Moreover, we use a Transformer[42] for text modality, and ViT[10] encoders for chart and sketch modalities, respectively. To achieve alignment across modalities within the joint embedding space, we implement a bidirectional hinge-based triplet ranking loss function and a two-level cross-entropy loss function. Through this process, we successfully align the three modalities and yield strong performance across four evaluation metrics surpassing a threshold of 0.85.
Visualization Referencing. In the visualization referencing module, we generate visualization references that comprehensively capture key data patterns inherent in an input table. We adopt a two-stage approach that first decomposes the input table $D$ into data facets, which are subsequently mapped to visualization references. In the facet generation stage, we employ a multifaceted process involving smoothing to preserve critical patterns while eliminating noise, as well as segmentation to extract meaningful and reliable patterns. Next in the visual mapping stage, these facets are mapped into fixed-size visualization references, effectively representing the patterns within the original tabular data.
Visualization Pruning. In the visualization pruning module, we introduce pruning and indexing techniques to enable timely visualization reference retrieval. Specifically, visualization pruning works by filtering out less informative visualization references among those with a high degree of similarity, thereby reducing the number of stored visualization references meanwhile keeping key data change patterns. In parallel, visualization indexing employs an open-source vector database, Chroma[1], to efficiently store and retrieve visualization references. Here, we use the fine-tuned LVLM to transform all the pruned visualization references into 512-dimensional vectors, which are then indexed by Chroma for rapid and accurate follow-up retrieval. This combination of pruning and indexing ensures VisLTR can swiftly manage and retrieve massive visualization references with precision.
Visualization Interaction. In the visualization interaction module, we introduce a “decomposing-execution-filling” strategy to enable the connection of multi-modal user queries with the stored embeddings of visualization references. Specifically, the user query is first decomposed into a visualization retrieval task. Second, the retrieval task is executed in Chrome to get a retrieved visualization reference. Afterward, with the help of the fine-tuned LVLM, the corresponding trend word of the retrieved visualization reference is generated and then filled into the textual response to the user. An interactive visual interface with multiple views and rich interactions is also developed to support the iterative table reasoning process.
4 Multi-modal Visualization Alignment
We set out to establish an LVLM that aligns visualization modalities of chart, sketch, and text in a single joint embedding space. This section presents the data preparation, including data augmentation and user labeling strategies, to address the issue of lacking training dataset (Sect. 4.1), followed by modality binding for establishing a joint embedding space that binds chart, sketch, and text modalities together (Sect. 4.2). In the end, we present the implementation details (Sect. 4.3).
Figure 2: An overview of visualization alignment module, which contains data preparation for chart-text and chart-sketch pairings, and achieves the alignment for chart, sketch, and text modalities.
4.1 Data Preparation
To train the multi-modal visualization alignment model, we need to first prepare a dataset containing user intentions, which is not covered by current open-sourced datasets. To achieve the alignment of sketches, charts, and text, we employ two strategies to construct the training data: (1) data augmentation for chart-text pairing, and (2) user labeling for chart-sketch pairing. Figure 2 uses an example to illustrate the two data preparation strategies.
(1) Data augmentation for chart-text pairing. To make up for the lack of data change pattern recognition in the existing table reasoning works, we should focus on the charts that describe data change patterns. We first survey existing chart-text datasets, and find that different textual contents are contained for different downstream tasks. We select the Chart-to-Text dataset[23] as the original dataset because it offers textual descriptions and underlying data tables for charts. Since we mainly focus on the data change patterns existing in charts, we filter the trend words from the textual descriptions. Initially, we manually extracted the contained trend words in five textual descriptions. These five descriptions with their corresponding trend words served as prompts for GPT-4 to perform an accurate trend word filtering process on the large Chart-to-Text dataset. In the filtering process, we find diversity in both trend words and charts. For trend words, we notice that different trend words in textual descriptions may indicate the same change pattern, such as ‘increasing’, ‘upward’, and ‘rising’, which typically indicate an upward trend in the data. We construct a dictionary of similar trend words, grouping the words describing similar patterns into the same category, and we get 23 categories ($N_{T}=23$) of trend words. For charts, we find that three chart types: line, bar, and area charts, are widely chosen to describe the data change patterns. To balance the number of three chart types contained in the final chart-text pairing dataset, we supplement each filtered chart with two other types of charts. In total, we create a total of 5,010 chart-text pairings.
Figure 3: The alignment pipeline for three modalities. Each modality is processed by a separate encoder, and the chart modality is set as the supervisor, guiding the alignment process.
(2) User labeling for chart-sketch pairing. When searching for hand-drawn sketch datasets associated with charts, we encounter a significant challenge: existing public sketch datasets do not adequately meet our need to convey users’ intentions about data change patterns. Specifically, the existing sketch datasets are not suitable for our work due to the following points. 1) Most open-sourced large human sketch datasets e.g.,[12], mainly focus on how users draw objects, rather than on visualizations. 2) LADV[33] and VisAtlas[55] are designed for chart type matching based on pixel-wise similarity, but not the contained data change patterns. 3) Sketch query systems, such as Qetch[35], focus on data-level similarities between hand-drawn sketches and the raw data, which lacks correspondence with fixed-size charts.
In order to collect chart-sketch pairings that conform to users’ intentions about data change patterns, we invite users to draw sketches and evaluate chart-sketch pairings. 1) Setup. Every user is given a training session to get familiar with experimental procedures and hand-drawn board operations. 2) Reference chart selection. Every user is given 100 real charts selected from our chart-text pairing dataset. The selected charts cover 23 categories of trend words, and contain bar, line, and area charts. 3) Drawing. The selected charts are displayed to the user one by one, and the user draws a sketch, which he/she thinks represents the change patterns in the displayed chart on the free-scale board. 4) Evaluation/Redraw. After drawing all the given charts, we show them the sketch with the corresponding chart one by one, and the user could delete or redraw the sketches they are not satisfied with. A total of 10 users (six women, four men) aged from 18 to 30 were invited. They all have knowledge of data visualization and are experienced in drawing sketches. In the end, we filter low-quality sketches, resulting in 1,214 pairs of chart-sketch pairings ready for training.
4.2 Modality Binding
We aim to bind all text $\mathbb{T}$, chart $\mathbb{C}$ and sketch $\mathbb{S}$ modalities together in a joint embedding space. Existing CLIP-based works predominantly focus on aligning pairs of modalities, while our objective is to accomplish the alignment of three modalities in the absence of the pairings between text and sketch modalities. As shown in Figure 3, we employ a two-stage strategy: first, we employ modality-specific encoders to transform each modality into embeddings of uniform length[51]. All three modalities are processed using the Transformer architecture, leveraging its self-attention mechanism to effectively capture complex dependencies within each modality.
Given a chart $C_{i}$ with its corresponding observation $T_{i}$ in text modality, and observation $S_{i}$ in sketch modality, we encode them into normalized embeddings: $p_{i}=f(C_{i})$, $q_{i}=f(S_{i})$, and $k_{i}=f(T_{i})$, where $f$ is a 63M-parameter 12-layer 512-wide Transformer with the architecture modifications with 8 attention heads[42] for text modality, and the Vision Transformer (ViT)[10] with a patch size of 32 for chart and sketch modalities. Each modality is processed by a separate encoder, which is designed to cater to the unique characteristics of the input. We add a modality-specific linear projection head on each encoder to obtain a fixed size $d$ dimensional embedding, which is normalized, and set as $p_{i}$, $q_{i}$ and $k_{i}$ for loss calculation. Besides facilitating learning, this setup allows us to initialize a subset of the encoders with weights from pre-trained CLIP[41] models. This initialization not only provides a strong starting point for our LVLM but also enables it to better understand and align the semantic content across different modalities.
During the second stage, we combine the two-level cross-entropy and a bidirectional hinge-based triplet ranking loss to guide the three modalities to align in the latent space. We elect the chart modality, $\mathbb{C}$, as the supervised modality because visualization references are in the form of charts, and the chart-text and chart-sketch pairings that span a wide range of change patterns are well collected. Inspired by VSE++[13], we address this by introducing the two-level cross-entropy and a bidirectional hinge-based triplet ranking loss, which makes the matched pairs have smaller entropies than unmatched ones. The formulation for the two cross-entropy functions is specified as:
$H_{1}(C,T)=-\sum(p_{i}\times\log(k_{i})),$
$H_{2}(C,S)=-\sum(p_{i}\times\log(q_{i})),$
where $H_{1}$ calculates the difference between charts $C$ and text $T$, and $(C,T)$ is the matched positive chart-text pair. $H_{2}$ calculates the difference between charts $C$ and sketches $S$, and $(C,S)$ is the matched positive chart-sketch pair.
We consider the loss function for the chart-text pairings as $L_{1}(C,T)$, and the loss function for the chart-sketch pairings as $L_{2}(C,S)$. Take chart and sketch modalities as an example, the loss is as follows:
$L_{2}(C,S)=[\alpha+H_{2}(C^{\prime},S)-H_{2}(C,S)]_{+}+[\alpha+H_{2}(C,S^{\prime})-H_{2}(C,S)]_{+},$
where $\alpha$ denotes the margin parameter, and $[x]_{+}\equiv\max(x,0)$. $C^{\prime}=\arg\max_{X\neq C}H_{2}(X,S)$ and $S^{\prime}=\arg\max_{Y\neq S}H_{2}(C,Y)$ denote the hardest negative corresponding to the positive pair $(C,S)$. The loss makes the chart embedding $p_{i}$ and the sketch embedding $q_{i}$ closer in the joint embedding space, and thus aligns $\mathbb{C}$ and $\mathbb{S}$. Set $L$ as the final loss of three modalities in the embedding space, that is, the sum of the two ranking losses:
$L(T,C,S)=L_{1}(C,T)+L_{2}(C,S).$
By minimizing both the loss between chart and text modalities, and the loss between chart and sketch modalities, we achieve the alignment between the similar embeddings of the three modalities in a joint embedding space.
4.3 Implementation Details
We took a pre-trained CLIP text-image model and fine-tuned it on the selected chart-text and chart-sketch pairings. In the training phases, we employed a strategy where we froze the ViT encoder and conducted fine-tuning for 30 epochs with a batch size of 32, and a learning rate of $1e^{-5}$. The model parameters were optimized using AdamW optimizer, employing Betas $(0.9,0.98)$, weight decay of 0.001, and incorporating of StepLR learning rate schedule. This static scheduler ensures that the learning rate gradually decreases during fine-tuning, facilitating convergence towards a local optimum. Through using the modalities paired with charts, i.e., pairs of the form $(\mathbb{C},\mathbb{T})$, and $(\mathbb{C},\mathbb{S})$ to align each the embeddings from text modality $\mathbb{T}$ and sketch modality $\mathbb{T}$ to those from charts. We observe that the pairs $(T,S)$ are aligned even though we only train using the pairs $(C,T)$ and $(C,M)$. This behavior is also validated by IMAGEBIND[16] so we reference to its statement that $(\mathbb{C},\mathbb{S})$ and $(\mathbb{C},\mathbb{T})$ are naturally aligned, and the text $\mathbb{T}$ and sketch modalities $\mathbb{S}$ constitute an emergency alignment.
5 VisLTR System
With the LVLM for multi-modal visualization alignment, we build VisLTR system that allows users to intuitively explore, comprehend, and interact with tabular data using visualizations as a proxy. This section details the implementation of visualization referencing (Sect. 5.1), visualization pruning (Sect. 5.2), and visualization interaction (Sect. 5.3) in VisLTR system.
5.1 Visualization Referencing
This module is designed to ensure that the generated visualization references can effectively capture comprehensive data patterns inherent in the raw tabular data. To this end, we design a two-stage approach that first generates data facets to capture all underlying data change patterns and then maps all data facets into visualization references.
Facet Generation. In the first stage, we aim to capture meaningful data facets of the input table as comprehensively as possible. If we use violent enumeration, it will produce a lot of candidate data facets, and many of them will not include data trends. To avoid the inefficient data facet generation, we adopt the idea of subset embedding [54] indicating that patterns of multi-dimensional data are often related to subsets rather than records. Specifically, we adopt a multifaced process that includes a dual strategy of smoothing and segmentation.
The multifaceted process we adopt incorporates a dual strategy focused on smoothing and segmentation. We use a time-series table as an instance for explanation, as it is rich in data change patterns that SQL-based LLM for table reasoning is not able to address. The initial step involves smoothing different variables to retain pivotal patterns while filtering out the noise. To achieve this, VisLTR employs Gaussian Smoothing, a technique favored for its property of maintaining shape consistency under the Fourier transform. Compared with widely used Moving Average (MA) or Weighted Moving Average (WMA)[18], Gaussian Smoothing can not only preserve key features while smoothing but also have flexibility in parameter control.
After smoothing, we focus on segmenting the table $D(d_{1},d_{2},\cdots,d_{n})$ into facets to extract meaningful and reliable patterns. In order to facilitate the discussion, we focus on the segmentation of one variable $d_{i}$ in the following. This is accomplished by identifying a set of data facets $r(d_{i})$ for each variable $d_{i}$, which represents a continuous subset of values. To ensure each data facet reflects a valid data pattern, we incorporate the Page-Hinckley Test (PHT)[38] to identify critical points that demote substantial changes. By setting these critical points as boundaries for the segments, we can delineate different facets of the variable $d_{i}$. The comprehensive set of data facets for the variable $d_{i}$ is represented by $r(d_{i})=PHT(d_{i})$.
Visual Mapping. The second stage is visual mapping, where we harness the matplotlib library to transform each data facet into a fixed-size visualization reference that is devoid of axes, specifically utilizing bar, line, and area charts to manifest the trend in each data facet. Furthermore, the uniformity provided by fixed-size visualization references improves the fine-tuned LVLM’s capacity to analyze and process them. The visualization references are represented as $C=(c_{1},\cdots,c_{n})$, where $c_{j}=vmap(r_{j}(d_{i}))$, where $r_{j}(d_{i})$ is a data facet selected from the set $r(d_{i})$, and $vmap$ represents the visual mapping process.
5.2 Visualization Pruning and Indexing
For a given table, we will generate a vast collection of visualization references, each of which is subsequently mapped into vector representations (i.e., visualization embeddings). The large number of visualization references poses significant challenges to efficient storage and retrieval. To address these challenges, we first introduce a visualization pruning technique to effectively reduce the overall number of visualization references, thereby reducing the required storage space and initially streamlining the retrieval process. However, even after pruning, the remaining number of visualization references may still be too large, demanding a visualization indexing technique to accelerate the retrieval process. The combined approach of visualization pruning and indexing, as illustrated in Figure 4, significantly improves VisLTR’s ability to manage and retrieve visualization references.
Visualization Pruning. The key idea of visualization pruning is to filter out less informative visualization references in advance, narrowing down the pool of candidates to those most likely to match the query. The main challenge is retaining highly valuable visualization references during pruning. To this end, we introduce a visualization pruning method to filter out less informative visualization references. For all visualization references for a variable $d_{i}$, it may encounter a scenario where two references exhibit similar change patterns, and one is entirely encompassed by the other. An example scenario is illustrated in Figure 4, where the ‘a peak’ change pattern exists in pink-boxed visualization references. In this case, the light-pink-boxed one has less information and is consequently filtered out. By leveraging the visualization pruning technique, we filtered out about two-thirds of the visualizations that contained less information, effectively reducing the storage footprint.
Visualization Indexing. For efficient retrieval of visualization references, we utilize an open-sourced vector database named Chroma[1], which enables effective indexing and retrieval of embeddings. Specifically, we first apply our fine-tuned LVLM to encode all visualization references into 512-dimensional vectors. Chroma automatically maintains an index document when encoding all visualization references for retrieval. These indexes compress vectors into codes of a fixed size, storing them in an array. For similarity measurement, we employ the dot product function. By leveraging Chroma’s advanced indexing capabilities, VisLTR efficiently speeds up the retrieval process to handle vast amounts of visualizations, ensuring both speed and accuracy.
Figure 4: The pipeline of visualization referencing and visualization pruning modules. Facing the massive generated visualization references, two techniques including, visualization pruning and indexing are used to accelerate the storage and retrieval process.
Figure 5: The interactive interface of VisLTR, which contains four main views and supports multi-modal interactions for table reasoning.
5.3 Visualization Interaction
The Decomposing-Execution-Filling Strategy. Inspired by Dater[56], we introduce a ‘decomposing-execution-filling’ strategy to decompose and parse multi-modal user queries and return results that meet user expectations. Specifically, we first decompose a user’s table reasoning queries and convert them into visualization reference retrieval operations, similar to text-to-SQL parsing (e.g.,[40, 48]). Afterward, we execute the visualization retrieval operations in Chroma and get the result for back-filling the final response.
Here, we employ two LLMs with in-context learning using prompts to decompose user queries and complete outputs respectively. Below we illustrate an example using the case shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A). Given the query “What is the price trend of Apple during March?”, we first decompose the question into a retrieval task “return the visualization reference of Apple price in March”, and a cross-modal alignment task “identify the trend in the retrieved visualization reference”. The corresponding visualization reference is first retrieved from Chroma, and then aligned with the ‘two peaks’ pattern to generate the final textual response “There is a two-peak price trend in Apple during March.”. Detailed prompts are provided in our open-sourced code.
Visual Interface. In addition to textual responses, VisLTR also incorporates a user-friendly interface to enable intuitive table reasoning and exploration. As shown in Figure 5, the interface includes four main views: (A) Data Table, (B) Main View, (C) Chat Box, and (D) Pattern View. When a user uploads a table, the back-end of VisLTR processes the tabular data into visualization references and stores them in the vector database, while the front-end interface displays the uploaded table in Data Table, and presents the visualization for the data in the Main View. By default, Main View uses an area graph for the overall context information and a line chart for the details. Various operations, such as selecting different variables or combinations of variables to display, are enabled in the drop-down box in the upper right corner of Main View. For the display of multi-variables, we use a multi-line chart with colors to encode different variables. The bottom area chart also allows for fixed-scale zooming and brushing, allowing users to observe detailed variable changes.
The Chat Box provides a dialogue window that supports multi-round conversations and multi-modal inputs including uploading charts, drawing sketches, and typing text. Textual responses to user queries are also displayed in the Chat Box. The responses are also connected to Main View, where the corresponding visualization reference is highlighted. Pattern View presents groups of typical data change patterns for the variable selected in Main View. These groups are ordered from top to bottom, with the most populated data change pattern at the top, taping to the least at the bottom. Users can select a specific group of data change patterns, and then the top-3 visualization references of this pattern will be overlayed by highlighting the background. If there is an overlap among the displayed visualization references, the overlapping area corresponds to a darker background color. For users wishing to delve into each reference in detail, users can click the ‘show’ button located in the upper right corner of the Main View to achieve a circular review of each one. Additionally, Pattern View provides multi-modal information (description text, different chart types, and candidate sketches) for a single variable’s typical change patterns, including descriptive text, various charts, and potential sketch candidates.
6 Evaluation
We conduct quantitative evaluations and case studies for a thorough evaluation of VisLTR. The quantitative evaluation experiments evaluate the alignment ability of our fine-tuned LVLM (Sect. 6.1). Next, we present two cases on actual application scenarios (Sects. 6.2 & 6.3), representing a possible pathway through multi-modal interactions to complete a table reasoning task.
6.1 Effectiveness of Multi-modal Visualization Alignment
We evaluate the alignment performance of our multi-modal visualization alignment model from two perspectives: quantitative evaluation on text-chart retrieval, and user recognition on sketch-chart retrieval.
6.1.1 Text-Chart Retrieval
To rigorously assess the performance of our fine-tuned visualization alignment model in aligning chart-text pairs, we ensure that the evaluation dataset includes chart types of bar, line, and area chart, as well as all $T_{N}=23$ categories of trend words. Specifically, we randomly select 20 percent of the data for each chart type and each trend category from the chart-text pairing dataset for evaluation.
Metrics. Considering the imbalance among different trend word categories, we set the following metrics for a comprehensive evaluation:
  • Top-1 Accuracy (Acc). The accuracy calculates the ratio of true accurately identified pairs to the total number of pairs examined.
  • Weighted Recall (WR). WR gauges the ability to correctly identify all relevant results among all actual positives. $WR=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}R_{i}$, where $R_{i}$ is the recall of the $i^{th}$ trend-word category.
  • Weighted Precision (WP). WP determines the proportion of true positives within all positively classified results. $WP=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}P_{i}$, where $P_{i}$ is the precision of the $i^{th}$ trend-word category.
  • Weighted F1-Score (WF). WF adopts the weighted average considering the unevenness of the categories, which is calculated using the formula: $WF=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}\frac{2\times P_{i}\times R_{i}}{(P_{i}+R_{i})}.$
Table 1: Experimental results on chart-to-text retrieval test for all and individual chart types.
Model All Bar Line Area
Ours Acc (%) 89.02 86.29 89.00 89.12
WR (%) 87.52 86.00 86.62 87.86
WP (%) 88.42 86.87 87.49 88.71
WF (%) 87.45 85.91 86.54 87.79
Results. The quantitative results are presented in Table 1. The overall results for all chart types confirm the robustness of our model in aligning chart and text modalities. Moreover, the results for each individual chart type also provide substantial evidence of the model’s effectiveness across various chart types. Our model has the effective accuracy, with performance metrics exceeding 85% across all chart types. Nevertheless, it is observed that bar charts present slightly lower metrics in comparison to other chart types. This variation may be attributed to the inherent characteristics of bar charts, which utilize separated bars to encode data, resulting in less distinctiveness compared to the line marks used in line charts and the area marks used in area charts. To enhance the accuracy for bar charts, factors such as bar widths need to be taken into consideration.
6.1.2 Sketch-Chart Retrieval
We validate the alignment for chart-sketch pairs through a user study. We selected three pre-trained models, CLIP[41], OpenCLIP[8], and UniChart[37] for comparison. Initially, we randomly selected 200 sketches from the chart-sketch dataset. Then, we tasked each of the four models with determining the closest chart in the intent space to each selected sketch, indicating the chart with the highest similarity. This process constructed 800 matched sketch-chart pairings for subsequent user study. We recruited 30 participants (15 women, and 15 men) with experience in building charts. Each participant is randomly given 20 sketches with their matched charts by each morel and asked to rate the similarity between pairs of matched sketches and charts for each model, based on their visual perception of the correspondences. The ratings are gathered using a 5-point Likert scale[22], which provides a quantitative measure of the user agreement or disagreement with the similarities of chart-sketch pairs.
As shown in Figure 6, our model achieved the highest similarity agreement among participants (Mean = 4.33, SD = 0.67), and the Unichart achieved the lowest agreement with Mean = 2.25 and SD = 1.17. Given the non-normal distribution of collected ratings, we initiated our analysis by conducting the Friedman test to evaluate the statistical significance of ratings among the four models. The p-value of the Friedman test is less than 0.01, indicating significant differences between models. To further pinpoint the specific differences among models, we employed the Wilcoxon signed-rank test for paired comparisons, accompanied by Bonferroni correction to control errors. The pairwise comparisons showed that our model significantly outperformed the others. The detailed comparison results are shown in Figure 6. The poor performance of UniChart may be due to its focus on OCR features rather than chart features, leading to a weaker grasp of the varied patterns present within charts and hand-drawn sketches. The comparison results underscore the importance of encoding patterns in our alignment model, confirm the judicious choice of CLIP-ViT as our base pre-trained model, and highlight the superiority of our fine-tuned model in binding sketch and chart modalities.
Figure 6: Ratings of the visual similarity between pairs of the matched sketches and charts across our fine-tuned LVLM and three pre-trained models based on a 5-point Likert scale.
6.2 Case Study 1: Financial Data Exploration
In this case, we consider a financial data exploration scenario. A market analyst, Mike, is studying the Dow Jones Index (DJI.)’s change patterns post-COVID-19. The used data measures the daily price movements of 30 large US companies, spanning from January 2nd, 2020 to December 29th, 2023. The data includes five columns: Date, Open, High, Low, and Close, detailing the trading metrics for each business day. Mike first loads the DJI. data into VisLTR and selects ‘Close’ as the displayed variable in the Main View. VisLTR initially generates 64,261 visualization references, among which 13,823 visualizations are pruned for exploration. The average retrieval time for querying a visualization reference stored in Chroma is within 0.6s.
As shown in Figure 7, Mike explores the data as follows.
  • Table Summarization. Mike first enters in the dialogue box “Describe the data.”. The response from VisLTR includes two parts, one is the basic table description “The table shows the market performance of the Dow Jones Index from January 2, 2020 to December 29, 2023, including the date, opening index, intraday high index, intraday low index, and closing index.”, and the other is the overall trend description “The data suffers a sharp drop and then a gradual recovery. There are lots of fluctuations during the overall upward trend.”.
  • Question & Answering. Given this description, Mike is interested in the “sharp drop” pattern, so he tries to locate the pattern by entering the text “Please give me more details about the mentioned ‘sharp drop’ pattern.”. Based on the context, VisLTR can easily locate the “sharp drop” pattern appears, and the matched interval is highlighted in Main View. At the same time, a text description “In March 2020, the Dow Jones Index experienced a ‘sharp drop’ pattern. This drop reflects significant market volatility amid growing concerns over the COVID-19 pandemic.” is returned in the dialogue box.
  • Visual Query. Mike then freely sketches a ‘w’ pattern [Uncaptioned image] in the empty canvas and enters “Are there patterns in the data similar to my hand-drawn sketch?”. VisLTR retrievals for the visualization references similar to the input hand-drawn sketch from the vector database, and returns the three most similar matches in Main View. To facilitate user understanding, VisLTR also returns a text description of the matched results: “The three intervals with similar patterns are from 8-Feb-21 to 16-Mar-21, 18-Feb-21 to 18-Mar-21, and 15-Dec-20 to 11-Jan-21. This pattern is recognized as a double-bottom pattern. It occurred in December 2020 and February 2021 suggesting that during both periods, the market demonstrated buyer interest at lower price levels after experiencing declines. This support prevented further downward movement. Higher prices prompted more buyers to participate, contributing to the Dow’s upward momentum.”.
In this way, Mike gets an overall understanding of the tabular data, and also typical trends and patterns within the data. The multi-modal interactions, together with visual interactions like selecting and zooming in the visual interface, enable Mike to easily understand the DJI. data and identify interesting data change patterns.
Figure 7: Case study 1: data exploration for DJI. form 2020 to 2023. Mike gets a complete understanding of the data via table summarization, question & answering, and visual query enabled by VisLTR.
Figure 8: Case study 2: multivariate analysis for the uploaded air pollutant dataset. With VisLTR, Emily identifies typical data change patterns and relationships between different air pollutants.
6.3 Case Study 2: Air Pollutant Analysis
In the second case, we demonstrate the workflow of mining relationships between data change patterns of different variables. A student, Emily, is seeking the correlations between different air pollutants. The used air pollutant dataset includes the daily air quality metrics from January 1, 2017 to December 31, 2018, which is collected by the Beijing US Embassy [^1] https://aqicn.org/city/beijing/us-embassy/cn/. The data includes variables of Date, PM25, O3, temperature, and relative humidity, which provides insights into the fluctuations in air quality and correlations between air pollutants. Emily first loads the pollutant data into VisLTR, which generates a multi-line chart with each line representing a variable. In the backend, VisLTR generates a total of 102,787 visualization references, of which 40,243 are stored in Chroma after pruning.
Figure 8 illustrates Emily’s analysis process using VisLTR to explore the multivariate data.
  • Free Exploration. To obtain the typical data change patterns of pollutants, Emily selects four pollutant indicators, PM25, O3, temperature, and relative humidity, and obtains their typical change patterns in the Pattern View. She notices that all the four indicators exhibit obvious daily fluctuations. Specifically, PM25 exhibits a ‘zigzag pattern with varying amplitudes.’, indicative of its dynamic nature and daily variability. Similarly, a typical pattern of O3 is ‘An irregular zigzag pattern with an overall fluctuating trend.’, further highlighting the pronounced daily jitters in pollutant levels. These observations motivate Emily to explore the relationship between these pollutant indicators over a long time interval.
[…]
In this manner, Emily gains a comprehensive insight into the multivariate air pollutant data, identifying both the typical data change patterns and correlations among various air pollutant indicators.
7 Discussion
The case studies have demonstrated VisLTR’s capability in identifying and explaining data change patterns within tabular data. By integrating visualization-in-the-loop, VisLTR not only complements SQL-based LLMs for table reasoning but also aligns with the exploratory data analysis (EDA) process. EDA often involves summarizing the main characteristics of a dataset using visualization, enabling users to understand the data’s structure, identify patterns, detect anomalies, and formulate hypotheses for further investigation. In this sense, the proposed framework not only streamlines the process of EDA but also expands its horizons, providing users with novel approaches to uncover hidden patterns within datasets.
Generalizability. Nevertheless, to fully realize the potential of the proposed framework, we need to enhance the generalizability of VisLTR from the following perspectives.
  • More table types. VisLTR is currently instantiated on time-series tables, which contain data change patterns that cannot be processed by existing SQL-based LLMs for table reasoning. VisLTR has the potential to be extended to more general table types by employing alternative data transformation methods within the visualization referencing module. For instance, VisLTR can be adapted to recognize and analyze the distribution patterns in a student grade table. In this scenario, a data aggregation method could be employed to identify grade distribution patterns, such as the normal distribution. Supporting a broader range of table types is feasible, leveraging methods such as subset embedding[54] and learning-based approaches[46].
  • General user queries. VisLTR focuses on tasks related to reasoning about data change patterns in tables. Therefore, we primarily provide text-based user queries related to data change patterns, such as “what type of change pattern exists in the data?” However, this narrow focus also limits our ability to handle other general user queries, including those related to causal reasoning, such as “why did Apple stock rise in March?” When posed with such questions, VisLTR may encounter the ‘hallucination’ issue, leading to incorrect queries and responses. To achieve a comprehensive exploration and reasoning of a data table, we require more general user queries, which may necessitate the incorporation of external knowledge sources.
  • Intent-enhanced sketches. Users’ intuition about sketching targets often emerges from the exploration process. As a reason, VisLTR offers insights into common patterns within the uploaded table. However, expressing complex and less common patterns solely through hand-drawn sketches can be challenging for users[27]. To address this challenge, several solutions can be considered. First, incorporating additional interactive features such as dragging, resizing, and erasing, sketches can provide users with greater flexibility in adjusting their intent[30]. Summarizing multi-attribute temporal patterns and introducing a more effective time-series segmentation method have shown promise in addressing this issue[44]. By implementing these methods, we can further enhance the capabilities of VisLTR to support sketches with enhanced intents.
Scaling up the dataset. The successful alignment of our model heavily relies on acquiring high-quality training data. Scaling up the dataset to enhance the generalizability of VisLTR presents significant challenges for the data augmentation and user labeling strategies currently employed. Collecting user-labeled sketches is time-consuming, and incorporating intents into sketches would further increase complexity. To address this limitation, one solution is to leverage crowdsourcing strategies, although this still requires significant manual effort. Another innovative technique would be to use large models to simulate and augment a variety of hand-drawn patterns. LLM-powered data augmentation has been a promising direction[9], and we are interested in using this approach for scaling up the dataset.
Potential scenarios. Besides table reasoning, the proposed framework holds potential for other application scenarios, such as safeguarding sensitive data that requires a balance between data privacy and utility. By transforming raw tabular data into visualizations, VisLTR shields the secure information contained within tables from users, while still allowing for examination and querying of the data patterns. Moreover, VisLTR could also find applications in fraud detection, as visualizations generated from transactional data could help identify suspicious patterns indicative of fraudulent activities.
8 Conclusion
This work introduces VisLTR system built upon a novel visualization-in-the-loop framework for table reasoning, serving as a complement to existing SQL-based table reasoning research. By utilizing visualizations as a proxy, the framework offers several advantages, including the ability to identify data change patterns, mitigate LLM drift, and support multi-modal interactions. The implementation of VisLTR requires an effective multi-modal visualization alignment model that binds chart, text, and sketch modalities together in a joint embedding space. This is accomplished by fine-tuning an LVLM with a new dataset of chart-text and chart-sketch pairings. Together with strategies employed for visualization referencing, indexing, and retrieval, we have created an interactive system that facilitates multi-modal interactions with tabular data. The feasibility of VisLTR is demonstrated through quantitative evaluations of the visualization alignment model and practical case studies. The proposed visualization-in-the-loop framework offers a new research direction for table reasoning. We aim to extend the support to a wider range of table types, accommodate more generalized user queries, and explore additional application scenarios in the future.
Acknowledgements.

### v2

@@
-VisLTR: Visualization-in-the-Loop Table Reasoning
+VisTR: Visualizations as Representations for Time-series Table Reasoning
 Abstract
-Table reasoning transforms user requirements into corresponding answers according to the provided table, which is often integrated with natural language interfaces for lay users to explore tabular data effortlessly.
-Recent research exploits large language models to facilitate table reasoning, by transforming vague user requirements into structured query languages (SQLs).
-However, these SQL-based approaches often overlook changes in data patterns, suffer from LLM drift, and limit exploration to only text queries.
-To this end, VisLTR is designed as a visualization-in-the-loop table reasoning framework that leverages visualizations as a proxy to provide concise data representations, capture interesting data patterns, and support cross-modal analysis.
-We describe VisLTR as a process consisting of four major modules: 1) visualization alignment that utilizes large vision-language models to align visualizations across various modalities, including chart, text, and sketch; 2) visualization referencing that decomposes a table into multifaceted visualization references that comprehensively represent the table; 3) visualization pruning that incorporates data and retrieval pruning to excise visualization references with poor information and enhance retrieval efficiency; and 4) visualization interaction that offers an interactive visual interface with multi-modal interactions for user-friendly table reasoning.
-Quantitative evaluation demonstrates the effectiveness of the alignment model in cross-modal visualization pairings.
-We further demonstrate applications of the framework on various table reasoning tasks such as table summarization and pattern detection.
+Table reasoning involves transforming natural language questions into corresponding answers based on the provided data table.
+Recent research exploits large language models (LLMs) to facilitate table reasoning, which however struggle with pattern recognition and lack support for visual-based pattern exploration.To address these limitations, we propose VisTR, a framework that leverages visualizations as representations to facilitate data pattern recognition and support cross-modal exploration.
+We describe VisTR as a process consisting of four major modules: 1) visualization alignment that utilizes multimodal LLMs to align visualizations across various modalities, including chart, text, and sketch; 2) visualization referencing that decomposes a table into multifaceted visualization references that comprehensively represent the table; 3) visualization pruning that incorporates data and retrieval pruning to excise visualization references with poor information and enhance retrieval efficiency; and 4) visualization interaction that offers an interactive visual interface with multimodal interactions for user-friendly table reasoning.
+Quantitative evaluation with existing multimodal LLMs demonstrates the effectiveness of the alignment model in cross-modal visualization pairings.
+We further illustrate the applicability of the proposed framework in various time-series table reasoning and exploration tasks.
 keywords:
-Table reasoning, visualization reference, multi-modal learning
+Table reasoning, visualization representation, multimodal LLMs
 […]
-VisLTR supports data change pattern identification and multi-modal explorations of text queries and hand-drawn sketches using visualizations as a proxy between a table and user requirements (bottom).
-The proposed approach advances existing SQL-based LLM methods for table reasoning that overlook changes in data patterns, suffer from LLM drift, and limit exploration to only text queries (top).
+VisTR leverages visualizations as the representations between a table and user intentions.
+The proposed framework advances existing LLM-based table reasoning methods by supporting data change pattern recognition (Q1), improving pattern recognition for long tables (Q2), and enabling visual-based pattern exploration (Q3).
 1 Introduction
-Table reasoning is the task of deriving answers from tables (i.e., tabular data) based on specific questions[58], supporting a wide range of applications such as fact verification[6] and question answering[21].
-The key challenge of table reasoning lies in understanding both the intent of queries (e.g., a text query) and the semantics of tables, and accurately generating answers that require logical reasoning.
-Existing methods for table reasoning can be broadly categorized into generic table reasoning that directly generate answers using pre-trained models or large language models (LLMs); and program-aided table reasoning that utilizes executable query languages (e.g., SQL or SPARQL) as the bridge between user queries and answers derived from the table[49].
-Generic table reasoning performs reasoning implicitly[49, 5], which, however, is poor at tracking how answers are derived and increases the risk of plausible but incorrect responses.
-Instead, program-aided table reasoning ensures precise data retrieval and offers transparency in the reasoning process, which is essential for applications such as fact verification[56, 7].
-Therefore, program-aided reasoning methods are attracting significant attention from both academia and industry[15].
-Despite the growing popularity of program-aided reasoning methods, there exist several limitations.
-One of the primary issues stems from SQL’s inherent lack of analytical functionality[24], which leads to an overlook of data change patterns.
-Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A) illustrates a scenario where users seek to examine the price trend of Apple during March.
-Despite successfully retrieving a sub-table containing the price of Apple in March, SQL-based LLMs fall short when it comes to further querying and discerning change patterns within this sub-table.
-Second, LLMs are known to encounter drift issues, which refer to the loss of information in the middle of long contexts.
-LLM drift is common when processing large data tables[4], where the LLM’s reasoning process may encounter errors due to the loss of information in the middle, as illustrated in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(B).
-Last, table reasoning is often integrated with natural language interfaces, to facilitate lay users’ exploration of tabular data effortlessly.
-However, there are scenarios in which laymen are unable to express their intentions in natural language but would prefer to use hand-drawn sketches.
-An example is shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(C), where a user may not know the term “two-valley” but can easily sketch this pattern.
-Existing SQL-based LLMs do not support non-verbal interactions, which are highly demanded.
-To overcome the limitations, we introduce a visualization-in-the-loop table reasoning framework that incorporates visualization references into large vision-language models (LVLMs) to complement SQL-based LLMs for table reasoning.
-Our key idea is to transform a tabular data into a set of meaningful and insightful visualizations, termed visualization references, which serve as a proxy to capture data insights and align with human intentions.
-This framework has several advantages compared to existing SQL-based table reasoning methods.
-First, visualization references can capture the short- and long-term data change patterns.
-Considering the price trend query as an example (Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A)), the retrieved visualization reference could capture the “two-peak” price trend of Apple during March.
-Second, regardless of the length of data facets, the encapsulated visualization references can capture and retain the data change patterns in fixed-size images.
-Adopting fixed-size visualization references as the representation of data facets can effectively circumvent the ‘drift’ issues encountered during reasoning.
-As Figure VisLTR: Visualization-in-the-Loop Table Reasoning(B) shows, the stock price change over a year can be represented as a single visualization, where the price bottom can be easily located in this reference.
-Moreover, leveraging visualizations as a proxy offers a new modality of hand-drawn sketches for interacting with the tabular data, which align more intuitively with human cognition[27].
-When querying for a hand-drawn sketch as shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(C), it is easier to retrieve a similar visualization reference among the candidates than to retrieve a similar change pattern in the tabular data[35].
-However, the strategy presents several challenges.
-First, data tables often contain a variety of long- and short-term patterns, which should be comprehensively captured by the visualization references.
-(C1) Ensuring the visualization references are comprehensive enough to accurately represent an input data table, thereby meeting the diverse analytical needs, poses the first challenge.
-Second, (C2) effectively aligning textual descriptions, visualization images, and hand-drawn sketches to bridge user queries with visualization references, meanwhile enabling multi-modal interactions presents a challenge.
-Existing models for vision-language tasks such as chart question answering (CQA)[19] are not specifically designed to meet this requirement, thus being less effective in identifying data change patterns.
+Table reasoning is a crucial natural language processing task that involves automatically extracting and inferring information from tables to answer specific questions[12, 70].
+This task supports a wide range of applications such as fact verification[6] and question answering[26], where each question-table pair constitutes an instance for reasoning[69].
+Given the superior commonsense and logical reasoning capabilities of large language models (LLMs)[63], utilizing them for table reasoning has emerged as the predominant approach[5, 12].
+Existing methods can be broadly categorized into generic table reasoning, where pre-trained models or LLMs directly generate answers, often enhanced by techniques such as designing prompts[5] to guide iterative reasoning or employing flexible tabular formats[57, 69]; and program-aided table reasoning that utilizes executable query languages (e.g., SQL, SPARQL, or Python) as the bridge between user queries and answers derived from the table[65, 9, 19].
+Despite their advancements, there exist several limitations.
+First, these methods struggle with pattern recognition, which involves the automatic identification and extraction of meaningful patterns within data, such as trends in time-series data[35, 8].
+For instance, Q1 in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning illustrates a scenario where users seek to examine the price trend of Apple during March.
+Existing LLM-based methods cannot correctly capture and respond to the data change patterns within the corresponding time period.
+The challenge becomes more pronounced when dealing with large-size tables, as LLMs are known to encounter ‘drift’ issues, where models lose information in the middle of long contexts[4].
+This loss of context can result in errors during the reasoning process, making it difficult to extract patterns accurately, as illustrated in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q2).
+Last, the methods lack support for visual-based pattern exploration, making it challenging for users to iteratively engage with data tables, particulalry for lay users.
+Rather than expressing their intentions precisely in natural language, laymen would prefer to use multimodal inputs like hand-drawn sketches or charts[41].
+An example is shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q3), where a user may not know the term “two-valley” but can easily recognize this visual pattern from the collected bar chart.
+However, existing LLM-based methods do not support such visual-based exploration, which are increasingly in demand.
+To overcome the limitations of existing LLM-based table reasoning methods, we introduce a novel framework that incorporates visualization as representation into multimodal LLMs (MLLMs).
+Our key idea is to transform data tables into a set of meaningful and insightful visualizations, termed visualization references, which serve as a concise data representation to capture data insights and align with human intentions.
+This framework offers several advantages that address the shortcomings of existing LLM-based table reasoning methods, effectively enhancing the table exploration process.
+First, visualization references can help recognize the short- and long-term data change patterns.
+For instance, when recognizing price trend in the given data table (Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q1)), the corresponding visualization reference could capture the “two-peak” price trend of Apple during March.
+Second, regardless of the length of data facets, the visualization references can capture and retain the data change patterns in fixed-size images.
+Adopting fixed-size visualization references as the representations of data facets can effectively circumvent the ‘drift’ issues encountered during reasoning.
+For Q2 in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning, the stock price change over a year can be represented as a single visualization, where the price bottom can be easily located in this reference.
+Moreover, leveraging visualizations as a representation offers new modalities for interacting with the tabular data, aligning more intuitively with human intentions[32].
+Combined with multimodal input, including text and chart (as shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q3)), it becomes easier to explore data tables than directly observing the raw table.
+We illustrate the effectiveness of VisTR on time-series data tables, which often feature large table sizes and complex data change patterns.
+In this context, the proposed framework also presents several challenges to be addressed.
+First, time-series data often contain a variety of long- and short-term patterns, which should be comprehensively captured by the visualization references.
+To ensure the visualization references are comprehensive enough to accurately represent a data table, thereby meeting the diverse analytical needs, poses the first challenge (C1).
+Second, to effectively align textual descriptions, visualization images, and hand-drawn sketches to bridge user intentions with visualization references, meanwhile enabling multimodal interactions presents a challenge (C2).
+Existing models for vision-language tasks such as chart question answering (CQA)[24, 68] are not specifically designed to meet this requirement, thus being less effective in identifying data change patterns.
 Last, to comprehensively represent a table, a large number of visualization references will be generated.
-(C3) Effectively indexing the visualization references and supporting efficient visualization retrieval for timely user response present a significant challenge.
-To evaluate the effectiveness of the proposed framework, we design a prototype system, which encompasses four main modules aimed at addressing the challenges C1--C3.
-The visualization referencing module decomposes an input table into data facets and processes the facets into corresponding visualization references that comprehensively capture the essence of the table.
+Effectively indexing the visualization references and supporting efficient visualization retrieval for timely user response present a significant challenge (C3).
+We design a prototype system that encompasses four main modules aimed at addressing the challenges C1--C3.
+First, the visualization referencing module decomposes an input table into data facets and processes the facets into corresponding visualization references that comprehensively capture the essence of the table.
 To ensure a comprehensive representation of data patterns in tabular data, we adopt a multifaceted process to get meaningful facets and then transform each facet into a fixed-size visualization reference (addressing C1).
-Second, in the visualization alignment module, we fine-tune a large vision-language model (LVLM) to connect visualizations of various modalities including charts, text, and sketches (addressing C2).
-To fine-tune the model, we construct a new cross-modal visualization dataset using data augmentation for chart-text pairing and user labeling for chart-sketch pairing, taking into account user intentions.
+Second, in the visualization alignment module, we fine-tune a MLLM to connect visualizations of various modalities including charts, text, and sketches (addressing C2).
+Specifically, we construct a new cross-modal visualization dataset using data augmentation for chart-text pairing and user labeling for chart-sketch pairing, taking into account user intentions.
 We further leverage the CLIP model[48] enriched with two-level cross-entropy loss and bidirectional hinge-based triplet ranking loss to enhance the alignment.
 Next, the visualization pruning module leverages visualization pruning to filter out less informative visualization references, and visualization indexing to enhance the efficiency of the retrieval process for the embeddings of visualization references stored in a vector database (addressing C3).
-Finally, we design a visualization interface that enables users’ exploration of time-series data tables through multi-modal interactions.
-Here, we incorporate a “decomposing-execution-filling” strategy to integrate the output generated by VisLTR with the input table.
-Specifically, a user query is initially decomposed into visualization retrieval tasks, which are executed in the vector database, and their corresponding text descriptions are extracted by the fine-tuned LVLM.
+Finally, we design a visualization interface that enables users’ exploration of time-series data tables through multimodal interactions.
+Here, we incorporate a “decomposing-execution-filling” strategy to integrate the output generated by VisTR with the input table.
+Specifically, the user input is initially decomposed into visualization retrieval tasks, which are executed in the vector database, and their corresponding text descriptions are extracted by the fine-tuned MLLM.
 Subsequently, these descriptions are filled back into the input table to generate textual responses for the user.
-Quantitative evaluations demonstrate the effectiveness of the fine-tuned LVLM, and case studies illustrate the system’s applicability in diverse application scenarios.
+Quantitative evaluations demonstrate the effectiveness of the fine-tuned MLLM in aligning different modalities.
+Additionally, case studies illustrate how the proposed framework addresses the limitations of existing LLM-based table reasoning methods through various time-series table reasoning and visual-based exploration tasks.
 Dataset, fine-tuned model, and evaluation results are released at https://osf.io/yun9d/.
 In summary, we make the following contributions:
-(1) Visualization-in-the-Loop Framework.
-We introduce a visualization-in-the-loop framework that leverages visualizations as the proxy to support table reasoning.
-The framework complements existing program-aided approaches for table reasoning, particularly in capturing data change patterns and facilitating non-verbal exploration.
-(2) Multi-modal Visualization Alignment.
-We design an effective LVLM that learns a joint embedding space across three visualization modalities - charts, text, and hand-drawn sketches.
-The model aligns multi-modal user intentions with visualization references, allowing users to explore and analyze tables through various interactions.
-(3) VisLTR System.
-We develop a prototype system that supports the exploration of large-scale and multiple time-series tables through intuitive user interactions.
-Various case studies and quantitative results prove the effectiveness and applicability of VisLTR.
+(1) Framework. We introduce a framework that leverages visualizations as the representations to enhance table reasoning, to facilitate data change pattern recognition and visual-based exploration.
+(2) Multimodal Visualization Alignment.
+We design an effective MLLM that learns a joint embedding space across three visualization modalities - charts, text, and hand-drawn sketches.
+The model aligns multimodal user intentions with visualization references, allowing users to explore and analyze tables through various interactions.
+(3) VisTR System.
+We develop a prototype system that supports the exploration and reasoning of large-scale and multiple time-series tables through intuitive user interactions.
+Various case studies and quantitative results prove the effectiveness and applicability of VisTR.
 2 Related Work
-2.1 Table Reasoning
-Table reasoning integrated with natural language (NL) interfaces supports lay users to explore tabular data effortlessly.
-This capability is essential for many complex tasks such as fact verification[6] and question answering[21].
-Existing table reasoning methods can be categorized into generic table reasoning and program-aided table reasoning[49].
+Table Reasoning.
+Table reasoning integrated with natural language (NL) interfaces enables lay users to explore tabular data effortlessly, supporting complex tasks such as fact verification[6] and question answering[26].
+Existing LLM-based table reasoning methods can be categorized into generic table reasoning and program-aided table reasoning[57].
 Generic reasoning methods employ LLMs to interpret and respond to free-form NL queries by generating answers directly from tables[5].
-Due to the complexity of table structure, these methods face difficulties in dealing with large tables.
-Moreover, explainability is limited due to their end-to-end nature.
-To facilitate complex table reasoning, program-aided table reasoning utilizes executable query languages, such as SQL, to form a more structured and accountable bridge between user quires and the derived answers[11, 31, 20].
-Recent research has mostly focused on leveraging LLMs to transform complex user utterances into SQL queries with data context awareness[56], using prompt engineering[25], chain-of-thought approaches[49, 57], and pre-training techniques[48].
-However, SQL-based LLM table reasoning approaches encounter intricate limitations due to their reliance on SQL.
+However, due to the complexity of table structure, these methods struggle with large tables and offer limited explainability because of their end-to-end nature.
+Program-aided table reasoning utilizes executable query languages, such as SQL, to form a structured and accountable bridge between user quires and the derived answers[14, 37, 25].
+Recent research has mostly focused on leveraging LLMs to transform complex user utterances into SQL queries with data context awareness[65], using prompt engineering[30], chain-of-thought[57, 66], and pre-training techniques[56].
+However, program-aided approaches encounter intricate limitations due to their reliance on query language.
 First, SQLs primarily cater to numerical queries from tables but lack support of analytical functionality[29], overlooking data change patterns.
 Second, when confronted with large tables, LLMs may encounter drift issues[4] while accessing information in the middle of long contexts.
-Moreover, in many scenarios, relying solely on textual inputs restricts the expression of non-verbal intentions, such as hand-drawn sketches.
-To overcome these limitations, this work introduces a novel framework that leverages visualizations as a proxy between tables and user queries, complementing SQL-based table reasoning by addressing challenges such as data change pattern identification and multi-modal interaction.
-The strategy is feasible, as visualizations can depict intuitive data patterns and illustrate complex underlying processes[14], facilitating swift and effective interpretation by the human brain[26].
+Moreover, in many scenarios, relying solely on textual inputs limits the expression of non-verbal intentions, such as hand-drawn sketches, reducing the flexibility of user interactions.
+To overcome these limitations, this work introduces a novel framework that leverages visualizations as a representation between tables and user intentions.
+Our inspiration is that visualizations can depict intuitive data patterns and illustrate complex underlying processes[18], aiding swift and effective interpretation by users[31].
 However, this approach also presents challenges to be tackled.
@@
 Second, there are no direct approaches available to connect free-form user queries with visualizations.
-Figure 1: An overview of the proposed framework, which mainly contains four main modules: 1) the visualization referencing module decomposes a raw data table into facets and maps them to visualization references; 2) the visualization pruning module leverages pruning and indexing techniques to ease the storage and improve the retrieve efficiency; 3) the visualization alignment module leverages a fine-tuned LVLM to enable alignment of various visualization modalities, including chart, text, and sketch; and 4) the visualization interaction module decomposes user queries into retrieval tasks with visualization references and utilizes the LVLM to generate the corresponding trend words to fill the textual responses for users.
-2.2 Multi-modal Visualization Alignment
-Given that visualizations are inherently stored as images, we opt to leverage emerging vision-language models to embed visualization images and facilitate multi-modal user interactions.
+Figure 1: An overview of the proposed framework, which mainly contains four main modules of visualization referencing, visualization pruning, visualization alignment, and visualization interaction.
+Multimodal Visualization Alignment.
+Given that visualizations are inherently stored as images, we opt to leverage emerging MLLMs to embed visualization images and facilitate multimodal user interactions.
 In recent years, significant advancements have been made in utilizing vision-language models for chart-based tasks, such as CQA[42, 43, 24], chart retrieval[60] and generation[50, 61].
 Though CQA models can provide answers for user utterances, these models only support text-chart interactions, lacking the capability for cross-modal interactions involving multiple modalities such as text, chart, and sketch preferred by laymen.
-Furthermore, these models are not specifically designed for effectively identifying data change patterns (see Sect.
+Furthermore, these models are not specifically designed for effectively recognizing data change patterns (see Sect.
 6.1), a gap that this work seeks to fill in the context of table reasoning.
 To this end, we choose to construct a novel vision-language model with an innovative visualization alignment strategy.
-Most existing multi-modal data alignment methods rely on contrastive language-image pre-training (CLIP)[41], which has demonstrated its adaptability and proficiency across various tasks[59, 29, 60], highlighting the effectiveness of joint embedding spaces.
+Most existing multimodal data alignment methods rely on contrastive language-image pre-training (CLIP)[48], which has demonstrated its adaptability and proficiency across various tasks[71, 34, 72], highlighting the effectiveness of joint embedding spaces.
 However, training an effective CLIP-based alignment model for visualizations is non-trivial.
@@
 Efforts have been made to construct datasets for promoting deep learning models for chart-text (e.g.,[28, 42]), chart-chart (e.g.,[40, 33, 64]), chart-sketch (e.g.,[41, 52, 38]) similarity measurement.
-However, these different modalities are not inherently unified in one model, and many of these datasets are not specifically designed for identifying data change patterns.
-To address this gap, we utilize data augmentation and user labeling strategies to supplement existing datasets in training the multi-modal visualization alignment model.
-2.3 Visualization Storage, Indexing, and Retrieval
+However, these different modalities are not inherently unified in one model, and many of these datasets are not specifically designed for recognizing data change patterns.
+To address this gap, we utilize data augmentation and user labeling strategies to supplement existing datasets in training the multimodal visualization alignment model.
+Visualization Storage, Indexing, and Retrieval.
 To comprehensively represent tabular data, our approach will generate a large number of visualization references and encode them into vector representations, thus raising new challenges for efficiently storing, indexing, and retrieving them.
@@
 In this work, we first introduce a visualization pruning method to filter out less informative visualizations, and effectively reduce storage demands while maintaining the quality of stored visualizations.
-We then integrate Chroma[1] into VisLTR and utilize an efficient graph-based index to store and index the large collection of visualization references.
-An Approximate k-Nearest Neighbors (ANN) algorithm is incorporated for quick and accurate retrieval of visualizations.
-The combination of pruning with indexing strategies allows for quick retrieval and handling timely user queries.
-3 The Visualization-in-the-Loop Framework
-The proposed visualization-in-the-loop framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Figure 1.
+We then integrate Chroma[1] into VisTR and utilize an efficient graph-based index to store and index the large collection of visualization references.
+An Approximate k-Nearest Neighbors (ANN) algorithm is incorporated to retrieve visualizations.
+The combination of pruning with indexing allows for quick retrieval and handling timely user queries.
+3 The Visualization Enhanced Framework
+The visualization enhanced framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Figure 1.
 Visualization Alignment.
-As the cornerstone of our framework, an LVLM is fine-tuned to establish a single joint embedding space for various modalities, including chart ($\mathbb{C}$), text ($\mathbb{T}$), and sketch ($\mathbb{S}$).
+As the cornerstone of our framework, a MLLM is fine-tuned to establish a single joint embedding space for various modalities, including chart ($\mathbb{C}$), text ($\mathbb{T}$), and sketch ($\mathbb{S}$).
 To address the issue of a lack of suitable training data, we first curate a new dataset with chart-text pairings augmented from existing Chart-to-Text dataset[28], and chart-sketch pairings through user labeling.
-We apply pre-trained CLIP[41] to fine-tune a multi-modal model on our composite dataset.
-Moreover, we use a Transformer[42] for text modality, and ViT[10] encoders for chart and sketch modalities, respectively.
+Utilizing the pre-trained CLIP[48], we fine-tune a multimodal LLM on our composite dataset, employing a Transformer[49] for text modality, and ViT[13] encoders for chart and sketch modalities, respectively.
 To achieve alignment across modalities within the joint embedding space, we implement a bidirectional hinge-based triplet ranking loss function and a two-level cross-entropy loss function.
 Through this process, we successfully align the three modalities and yield strong performance across four evaluation metrics surpassing a threshold of 0.85.
+A detailed description of the module is provided in Sect.
+4.
 Visualization Referencing.
 In the visualization referencing module, we generate visualization references that comprehensively capture key data patterns inherent in an input table.
-We adopt a two-stage approach that first decomposes the input table $D$ into data facets, which are subsequently mapped to visualization references.
-In the facet generation stage, we employ a multifaceted process involving smoothing to preserve critical patterns while eliminating noise, as well as segmentation to extract meaningful and reliable patterns.
-Next in the visual mapping stage, these facets are mapped into fixed-size visualization references, effectively representing the patterns within the original tabular data.
+This module employs a two-stage approach that first decomposes the input table $D$ into data facets, which are then mapped into fixed-size visualization references.
+The description of this module is detailed in Sect.
+5.1.
 Visualization Pruning.
 In the visualization pruning module, we introduce pruning and indexing techniques to enable timely visualization reference retrieval.
-Specifically, visualization pruning works by filtering out less informative visualization references among those with a high degree of similarity, thereby reducing the number of stored visualization references meanwhile keeping key data change patterns.
-In parallel, visualization indexing employs an open-source vector database, Chroma[1], to efficiently store and retrieve visualization references.
-Here, we use the fine-tuned LVLM to transform all the pruned visualization references into 512-dimensional vectors, which are then indexed by Chroma for rapid and accurate follow-up retrieval.
-This combination of pruning and indexing ensures VisLTR can swiftly manage and retrieve massive visualization references with precision.
+Specifically, visualization pruning works by filtering out less informative visualization references by measuring the similarity between visualization vectors using Euclidean distance and applying a pruning threshold.
+This process reduces the number of stored references while preserving essential data change patterns.
+Concurrently, visualization indexing leverages an open-source vector database, Chroma[1], to efficiently store and retrieve visualization references.
+Here, we use the fine-tuned MLLM to transform all the pruned visualization references into 512-dimensional vectors, which are then indexed by Chroma for rapid and accurate follow-up retrieval.
+This combination of pruning and indexing ensures VisTR can swiftly manage and retrieve massive visualization references with precision.
+This module is detailed in Sect.
+5.2.
 Visualization Interaction.
-In the visualization interaction module, we introduce a “decomposing-execution-filling” strategy to enable the connection of multi-modal user queries with the stored embeddings of visualization references.
+In the visualization interaction module, we introduce a “decomposing-execution-filling” strategy to enable the connection of multimodal user queries with the stored embeddings of visualization references.
 Specifically, the user query is first decomposed into a visualization retrieval task.
 Second, the retrieval task is executed in Chrome to get a retrieved visualization reference.
-Afterward, with the help of the fine-tuned LVLM, the corresponding trend word of the retrieved visualization reference is generated and then filled into the textual response to the user.
-An interactive visual interface with multiple views and rich interactions is also developed to support the iterative table reasoning process.
-4 Multi-modal Visualization Alignment
-We set out to establish an LVLM that aligns visualization modalities of chart, sketch, and text in a single joint embedding space.
+Afterward, with the help of the fine-tuned MLLM, the corresponding trend word of the retrieved visualization reference is generated and then filled into the textual response to the user.
+An interactive visual interface with multiple views and rich interactions is also developed to support the iterative table reasoning process, transitioning from abstract to detailed exploration, as detailed in Sect.
+5.3.
+4 Multimodal Visualization Alignment
+We set out to establish a MLLM that aligns visualization modalities of chart, sketch, and text in a single joint embedding space.
 This section presents the data preparation, including data augmentation and user labeling strategies, to address the issue of lacking training dataset (Sect.
@@
 4.1 Data Preparation
-To train the multi-modal visualization alignment model, we need to first prepare a dataset containing user intentions, which is not covered by current open-sourced datasets.
+To train the multimodal visualization alignment model, we need to first prepare a dataset containing user intentions, which is not covered by current open-sourced datasets.
 To achieve the alignment of sketches, charts, and text, we employ two strategies to construct the training data: (1) data augmentation for chart-text pairing, and (2) user labeling for chart-sketch pairing.
@@
 In total, we create a total of 5,010 chart-text pairings.
-Figure 3: The alignment pipeline for three modalities. Each modality is processed by a separate encoder, and the chart modality is set as the supervisor, guiding the alignment process.
 (2) User labeling for chart-sketch pairing.
@@
 In the end, we filter low-quality sketches, resulting in 1,214 pairs of chart-sketch pairings ready for training.
+Figure 3: The alignment pipeline for three modalities. Each modality is processed by a separate encoder, and the chart modality is set as the supervisor, guiding the alignment process.
 4.2 Modality Binding
@@
 Each modality is processed by a separate encoder, which is designed to cater to the unique characteristics of the input.
-We add a modality-specific linear projection head on each encoder to obtain a fixed size $d$ dimensional embedding, which is normalized, and set as $p_{i}$, $q_{i}$ and $k_{i}$ for loss calculation.
+We add a modality-specific linear projection head on each encoder to obtain a fixed size $d$ dimensional embedding, which is normalized, and set as $p_{i}$, $q_{i}$, and $k_{i}$ for loss calculation.
 Besides facilitating learning, this setup allows us to initialize a subset of the encoders with weights from pre-trained CLIP[48] models.
-This initialization not only provides a strong starting point for our LVLM but also enables it to better understand and align the semantic content across different modalities.
+This initialization not only provides a strong starting point for our MLLM but also enables it to better understand and align the semantic content across different modalities.
 During the second stage, we combine the two-level cross-entropy and a bidirectional hinge-based triplet ranking loss to guide the three modalities to align in the latent space.
@@
 The loss makes the chart embedding $p_{i}$ and the sketch embedding $q_{i}$ closer in the joint embedding space, and thus aligns $\mathbb{C}$ and $\mathbb{S}$.
-Set $L$ as the final loss of three modalities in the embedding space, that is, the sum of the two ranking losses:
-$L(T,C,S)=L_{1}(C,T)+L_{2}(C,S).$
-By minimizing both the loss between chart and text modalities, and the loss between chart and sketch modalities, we achieve the alignment between the similar embeddings of the three modalities in a joint embedding space.
+Set $L$ as the final loss of three modalities in the embedding space, that is, the sum of the two ranking losses: $L(T,C,S)=L_{1}(C,T)+L_{2}(C,S).$ By minimizing both the loss between chart and text modalities, and the loss between chart and sketch modalities, we achieve the alignment between the similar embeddings of the three modalities in a joint embedding space.
 4.3 Implementation Details
@@
 This behavior is also validated by IMAGEBIND[20] so we reference to its statement that $(\mathbb{C},\mathbb{S})$ and $(\mathbb{C},\mathbb{T})$ are naturally aligned, and the text $\mathbb{T}$ and sketch modalities $\mathbb{S}$ constitute an emergency alignment.
-5 VisLTR System
-With the LVLM for multi-modal visualization alignment, we build VisLTR system that allows users to intuitively explore, comprehend, and interact with tabular data using visualizations as a proxy.
+5 VisTR System
+With the MLLM for multimodal visualization alignment, we build VisTR system that allows users to intuitively explore, comprehend, and interact with tabular data using visualizations as the representation.
 This section details the implementation of visualization referencing (Sect.
@@
 5.2), and visualization interaction (Sect.
-5.3) in VisLTR system.
+5.3) in VisTR system.
 5.1 Visualization Referencing
@@
 To avoid the inefficient data facet generation, we adopt the idea of subset embedding [62] indicating that patterns of multi-dimensional data are often related to subsets rather than records.
-Specifically, we adopt a multifaced process that includes a dual strategy of smoothing and segmentation.
-The multifaceted process we adopt incorporates a dual strategy focused on smoothing and segmentation.
-We use a time-series table as an instance for explanation, as it is rich in data change patterns that SQL-based LLM for table reasoning is not able to address.
-The initial step involves smoothing different variables to retain pivotal patterns while filtering out the noise.
-To achieve this, VisLTR employs Gaussian Smoothing, a technique favored for its property of maintaining shape consistency under the Fourier transform.
+Specifically, we adopt a multifaced process that incorporates a dual strategy of smoothing and segmentation.
+Consider an input time-series table $D(timestamp,d_{1},d_{2},\cdots,d_{n})$, where each $d_{i}$ represents a distinct variable containing a sequence of data points indexed by the timestamp.
+The first step is to smooth each variable $d_{i}$ to retain essential patterns while filtering out the noise.
+To achieve this, VisTR employs Gaussian Smoothing, a technique favored for maintaining shape consistency under the Fourier transform[53].
+Formally, for each each variable $d_{i}$, the smoothed series $\widetilde{d_{i}}$ is obtained as: $\widetilde{d_{i}}=GaussianSmoothing(d_{i})$.
 Compared with widely used Moving Average (MA) or Weighted Moving Average (WMA)[23], Gaussian Smoothing can not only preserve key features while smoothing but also have flexibility in parameter control.
-After smoothing, we focus on segmenting the table $D(d_{1},d_{2},\cdots,d_{n})$ into facets to extract meaningful and reliable patterns.
-In order to facilitate the discussion, we focus on the segmentation of one variable $d_{i}$ in the following.
-This is accomplished by identifying a set of data facets $r(d_{i})$ for each variable $d_{i}$, which represents a continuous subset of values.
-To ensure each data facet reflects a valid data pattern, we incorporate the Page-Hinckley Test (PHT)[38] to identify critical points that demote substantial changes.
-By setting these critical points as boundaries for the segments, we can delineate different facets of the variable $d_{i}$.
-The comprehensive set of data facets for the variable $d_{i}$ is represented by $r(d_{i})=PHT(d_{i})$.
+After smoothing, we segment each smoothed variable $\widetilde{d_{i}}$ into distinct facets to extract meaningful and reliable patterns.
+The segmentation process is defined as $R(\widetilde{d_{i}})=\{r_{1}(\widetilde{d_{i}}),r_{2}(\widetilde{d_{i}}),\cdots r_{m}(\widetilde{d_{i}})\}$, where $r_{j}(\widetilde{d_{i}})$ represents a subset (termed as a facet) of $\widetilde{d_{i}}$.
+To ensure each data facet reflects a valid data pattern, we incorporate the Page-Hinckley Test (PHT)[44] to detect critical points where shifts occur, thereby delineating the boundaries between different facets.
+Specifically, the segmentation process can be formalized as $R(\widetilde{d_{i}})=PHT(\widetilde{d_{i}})$.
+Additionally, we aggregate continuous subsets to form new facets, enabling the inclusion of segments with varying lengths.
 Visual Mapping.
-The second stage is visual mapping, where we harness the matplotlib library to transform each data facet into a fixed-size visualization reference that is devoid of axes, specifically utilizing bar, line, and area charts to manifest the trend in each data facet.
-Furthermore, the uniformity provided by fixed-size visualization references improves the fine-tuned LVLM’s capacity to analyze and process them.
-The visualization references are represented as $C=(c_{1},\cdots,c_{n})$, where $c_{j}=vmap(r_{j}(d_{i}))$, where $r_{j}(d_{i})$ is a data facet selected from the set $r(d_{i})$, and $vmap$ represents the visual mapping process.
+The second stage is visual mapping, where we transform each data facet into a standard visual representation.
+This process leverages the matplotlib library to generate fixed-size visualization references devoid of axes, specifically utilizing bar, line, and area charts to manifest the trend within each data facet.
+The uniformity provided by fixed-size visualization references improves the fine-tuned MLLM’s capacity to analyze and process them.
+For each data facet $r_{j}(\widetilde{d_{i}})\in R(\widetilde{d_{i}})$, the visual mapping process transforms $r_{j}(\widetilde{d_{i}})$ into a fixed-size visualization reference $c_{j}$, which is formalized as $c_{j}=vmap(r_{j}(\widetilde{d_{i}}))$.
+The collection of all visualization references derived from the data facets of a particular variable $d_{i}$ is represented as $C(d_{i})=\{c_{1},c_{2},\cdots,c_{m}\}$.
+Collectively, for all variables in the input table $D$, the visualization references are denoted as $C=\bigcup_{i=1}^{n}C(d_{i})=\{c_{i,j}|d_{i}\in D,r_{j}(d_{i})\in R(d_{i})\}$.
 5.2 Visualization Pruning and Indexing
-For a given table, we will generate a vast collection of visualization references, each of which is subsequently mapped into vector representations (i.e., visualization embeddings).
+For a given table, we will generate a vast collection of visualization references, each of which is subsequently mapped into vector representations (i.e., embeddings).
 The large number of visualization references poses significant challenges to efficient storage and retrieval.
-To address these challenges, we first introduce a visualization pruning technique to effectively reduce the overall number of visualization references, thereby reducing the required storage space and initially streamlining the retrieval process.
-However, even after pruning, the remaining number of visualization references may still be too large, demanding a visualization indexing technique to accelerate the retrieval process.
-The combined approach of visualization pruning and indexing, as illustrated in Figure 4, significantly improves VisLTR’s ability to manage and retrieve visualization references.
+To address these challenges, we design a combined approach of visualization pruning and indexing, as illustrated in Figure 4, to improve VisTR’s ability to manage and retrieve visualization references.
+Figure 4: The pipeline of visualization referencing and visualization pruning modules. Visualization pruning and indexing are used to accelerate the storage and retrieval process.
 Visualization Pruning.
-The key idea of visualization pruning is to filter out less informative visualization references in advance, narrowing down the pool of candidates to those most likely to match the query.
-The main challenge is retaining highly valuable visualization references during pruning.
-To this end, we introduce a visualization pruning method to filter out less informative visualization references.
+Here, the key idea is to filter out less informative visualization references in advance while retaining those that provide significant insights.
+The main challenge lies in discerning and preserving highly valuable visualization references during pruning.
 For all visualization references for a variable $d_{i}$, it may encounter a scenario where two references exhibit similar change patterns, and one is entirely encompassed by the other.
+We quantify the similarity between visualization vectors using the Euclidean distance $d(c_{j},c_{k})$ between the two vectors $c_{j}$ and $c_{k}$.
+We introduce a pruning threshold to determine whether two visualization vectors are sufficiently similar to warrant pruning.
+If $d(c_{j},c_{k})$ is less than the set pruning threshold, the vector covering a shorter time range is filtered out.
 An example scenario is illustrated in Figure 4, where the ‘a peak’ change pattern exists in pink-boxed visualization references.
 In this case, the light-pink-boxed one has less information and is consequently filtered out.
-By leveraging the visualization pruning technique, we filtered out about two-thirds of the visualizations that contained less information, effectively reducing the storage footprint.
+Figure 5: The relationship between the number of visualization references, table rows, and the pruning threshold.
+The remaining visualization references are close to approximately 20% of the raw number when the pruning threshold is set to 1.
+The pruning threshold acts as a predefined parameter that affects the number of stored visualization references.
+To assess its impact, we designed an experiment with four values: 0.5, 0.8, 1, and 1.5, and applied the thresholds to tables in a range of row numbers.
+Specifically, the row numbers were considered based on the typical sizes used in existing LLM-based table reasoning methods, which typically support data tables with fewer than 50 rows.
+For instance, Tabfact[7], a commonly used dataset containing over 16k Wiki tables, has an average of 12.96 rows per table (max.=48, min.=1, std.=8.46).
+Given the focus on time-series data tables and long-term trend analysis, we extended the experiment to tables with up to 1000 rows, approximately encompassing four trading years.
+We calculated the average number of visualization references across three sub-tables with corresponding table sizes extracted from the DJI.
+data table for each pruning threshold.
+Figure 5 shows the relationship between the number of stored visualization references, the number of data table rows, and the pruning thresholds.
+When the pruning threshold is set to 1, the number of stored visualization references approximated 20% of the raw reference number.
+This aligns with the well-known Pareto principle (80/20 rule)[45], which is widely leveraged in various optimization strategies to balance efficiency and information retention.
+As such, we also set the pruning threshold to be 1.
+This strikes a balance between maintaining a rich representation set and reducing the storage footprint.
 Visualization Indexing.
 For efficient retrieval of visualization references, we utilize an open-sourced vector database named Chroma[1], which enables effective indexing and retrieval of embeddings.
-Specifically, we first apply our fine-tuned LVLM to encode all visualization references into 512-dimensional vectors.
+Specifically, we first apply our fine-tuned MLLM to encode all visualization references into 512-dimensional vectors.
 Chroma automatically maintains an index document when encoding all visualization references for retrieval.
@@
 For similarity measurement, we employ the dot product function.
-By leveraging Chroma’s advanced indexing capabilities, VisLTR efficiently speeds up the retrieval process to handle vast amounts of visualizations, ensuring both speed and accuracy.
-Figure 4: The pipeline of visualization referencing and visualization pruning modules.
-Facing the massive generated visualization references, two techniques including, visualization pruning and indexing are used to accelerate the storage and retrieval process.
-Figure 5: The interactive interface of VisLTR, which contains four main views and supports multi-modal interactions for table reasoning.
+By leveraging Chroma’s advanced indexing capabilities, VisTR efficiently speeds up the retrieval process to handle vast amounts of visualizations, ensuring both speed and accuracy.
+Figure 6: The interactive interface of VisTR, which contains four main views and supports multimodal interactions for table reasoning.
 5.3 Visualization Interaction
 The Decomposing-Execution-Filling Strategy.
-Inspired by Dater[56], we introduce a ‘decomposing-execution-filling’ strategy to decompose and parse multi-modal user queries and return results that meet user expectations.
+Inspired by Dater[65], we introduce a ‘decomposing-execution-filling’ strategy to decompose and parse multimodal user queries and return results that meet user expectations.
 Specifically, we first decompose a user’s table reasoning queries and convert them into visualization reference retrieval operations, similar to text-to-SQL parsing (e.g.,[47, 56]).
@@
 Here, we employ two LLMs with in-context learning using prompts to decompose user queries and complete outputs respectively.
-Below we illustrate an example using the case shown in Figure VisLTR: Visualization-in-the-Loop Table Reasoning(A).
-Given the query “What is the price trend of Apple during March?”, we first decompose the question into a retrieval task “return the visualization reference of Apple price in March”, and a cross-modal alignment task “identify the trend in the retrieved visualization reference”.
+Below we illustrate an example using the case shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(A).
+Given the query “What is the price trend of Apple during March?”, we first decompose the question into a retrieval task “return the visualization reference of Apple price in March”, and a cross-modal alignment task “recognize the trend in the retrieved visualization reference”.
 The corresponding visualization reference is first retrieved from Chroma, and then aligned with the ‘two peaks’ pattern to generate the final textual response “There is a two-peak price trend in Apple during March.”.
@@
 Visual Interface.
-In addition to textual responses, VisLTR also incorporates a user-friendly interface to enable intuitive table reasoning and exploration.
+In addition to textual responses, VisTR also incorporates a user-friendly interface to enable intuitive table reasoning and exploration.
 As shown in Figure 6, the interface includes four main views: (A) Data Table, (B) Main View, (C) Chat Box, and (D) Pattern View.
-When a user uploads a table, the back-end of VisLTR processes the tabular data into visualization references and stores them in the vector database, while the front-end interface displays the uploaded table in Data Table, and presents the visualization for the data in the Main View.
+When a user uploads a table, the back-end of VisTR processes the tabular data into visualization references and stores them in the vector database, while the front-end interface displays the uploaded table in Data Table, and presents the visualization for the data in the Main View.
 By default, Main View uses an area graph for the overall context information and a line chart for the details.
@@
 The bottom area chart also allows for fixed-scale zooming and brushing, allowing users to observe detailed variable changes.
-The Chat Box provides a dialogue window that supports multi-round conversations and multi-modal inputs including uploading charts, drawing sketches, and typing text.
+The Chat Box provides a dialogue window that supports multi-round conversations and multimodal inputs including uploading charts, drawing sketches, and typing text.
 Textual responses to user queries are also displayed in the Chat Box.
@@
 For users wishing to delve into each reference in detail, users can click the ‘show’ button located in the upper right corner of the Main View to achieve a circular review of each one.
-Additionally, Pattern View provides multi-modal information (description text, different chart types, and candidate sketches) for a single variable’s typical change patterns, including descriptive text, various charts, and potential sketch candidates.
+Additionally, Pattern View provides multimodal information (description text, different chart types, and candidate sketches) for a single variable’s typical change patterns, including descriptive text, various charts, and potential sketch candidates.
 6 Evaluation
-We conduct quantitative evaluations and case studies for a thorough evaluation of VisLTR.
-The quantitative evaluation experiments evaluate the alignment ability of our fine-tuned LVLM (Sect.
-6.1).
-Next, we present two cases on actual application scenarios (Sects.
-6.2 & 6.3), representing a possible pathway through multi-modal interactions to complete a table reasoning task.
-6.1 Effectiveness of Multi-modal Visualization Alignment
-We evaluate the alignment performance of our multi-modal visualization alignment model from two perspectives: quantitative evaluation on text-chart retrieval, and user recognition on sketch-chart retrieval.
+To thoroughly evaluate VisTR, we conduct both quantitative evaluations and case studies.
+In the quantitative evaluations (Sect.
+6.1), we assess the alignment ability of our fine-tuned MLLM.
+Subsequently, we present two cases on actual application scenarios (Sects.
+6.2 & 6.3), representing a possible pathway through multimodal interactions to complete a time-series table reasoning task.
+6.1 Effectiveness of Multimodal Visualization Alignment
+We evaluate the alignment performance of our multimodal visualization alignment model from two perspectives: quantitative evaluation on text-chart retrieval, and user recognition on sketch-chart retrieval.
+6.1.1 Baselines and Metrics
+We compare our fine-tuned MLLM against several pre-trained models: CLIP[48], OpenCLIP[10], and UniChart[43].
+Our MLLM is a fine-tuned CLIP model on the composite dataset.
+To highlight the necessity and impact of the fine-tuning process, we compare it with the zero-shot CLIP and OpenCLIP models as an ablation study.
+Additionally, we validate the effectiveness of our fine-tuned CLIP by comparing it against OpenCLIP and Unichart, confirming the appropriateness of our model selection.
+VisTR enables comprehensive multimodal binding among different modalities, rather than performing direct subsequence or sketch matching.
+In this sense, we omit the comparison with existing sketch query[41, 17] or sequence matching systems[21, 67].
+Considering the imbalance among different trend word categories, we set two metrics for a comprehensive evaluation:
+  • Top-1 Accuracy (Acc). The accuracy calculates the ratio of true accurately identified pairs to the total number of pairs examined.
+  • Weighted F1-Score (WF).
+WF adopts the weighted average considering the unevenness of the categories, which is calculated as: $WF=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}\frac{2\times P_{i}\times R_{i}}{(P_{i}+R_{i})}$, where $R_{i}$ and $P_{i}$ are the recall and precision of the $i^{th}$ trend-word category, respectively.
 6.1.2 Text-Chart Retrieval
 To rigorously assess the performance of our fine-tuned visualization alignment model in aligning chart-text pairs, we ensure that the evaluation dataset includes chart types of bar, line, and area chart, as well as all $T_{N}=23$ categories of trend words.
-Specifically, we randomly select 20 percent of the data for each chart type and each trend category from the chart-text pairing dataset for evaluation.
-Metrics. Considering the imbalance among different trend word categories, we set the following metrics for a comprehensive evaluation:
-  • Top-1 Accuracy (Acc). The accuracy calculates the ratio of true accurately identified pairs to the total number of pairs examined.
-  • Weighted Recall (WR).
-WR gauges the ability to correctly identify all relevant results among all actual positives.
-$WR=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}R_{i}$, where $R_{i}$ is the recall of the $i^{th}$ trend-word category.
-  • Weighted Precision (WP).
-WP determines the proportion of true positives within all positively classified results.
-$WP=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}P_{i}$, where $P_{i}$ is the precision of the $i^{th}$ trend-word category.
-  • Weighted F1-Score (WF).
-WF adopts the weighted average considering the unevenness of the categories, which is calculated using the formula: $WF=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}\frac{2\times P_{i}\times R_{i}}{(P_{i}+R_{i})}.$
+Specifically, we randomly select 20% of the data for each chart type and each trend category from the chart-text pairing dataset for evaluation.
 Table 1: Experimental results on chart-to-text retrieval test for all and individual chart types.
 Model All Bar Line Area
-Ours Acc (%) 89.02 86.29 89.00 89.12
-WR (%) 87.52 86.00 86.62 87.86
-WP (%) 88.42 86.87 87.49 88.71
+Ours (Fine-tuned CLIP) Acc (%) 89.02 86.29 89.00 89.12
 WF (%) 87.45 85.91 86.54 87.79
+Fine-tuned OpenCLIP Acc (%) 50.74 50.36 51.86 50.24
+WF (%) 49.77 48.04 46.41 49.73
+Zero-shot CLIP Acc (%) 0.82 1.10 0.72 1.02
+WF (%) 0.36 0.29 0.11 0.07
+Zero-shot OpenCLIP Acc (%) 7.71 8.20 5.27 1.08
+WF (%) 5.37 3.78 1.95 0.30
 Results.
 The quantitative results are presented in Table 1.
-The overall results for all chart types confirm the robustness of our model in aligning chart and text modalities.
-Moreover, the results for each individual chart type also provide substantial evidence of the model’s effectiveness across various chart types.
-Our model has the effective accuracy, with performance metrics exceeding 85% across all chart types.
-Nevertheless, it is observed that bar charts present slightly lower metrics in comparison to other chart types.
-This variation may be attributed to the inherent characteristics of bar charts, which utilize separated bars to encode data, resulting in less distinctiveness compared to the line marks used in line charts and the area marks used in area charts.
-To enhance the accuracy for bar charts, factors such as bar widths need to be taken into consideration.
+Our fine-tuned MLLM demonstrates robust performance in aligning chart and text modalities, with overall accuracy and weighted F1-score exceeding 85% across all chart types.
+The results for each individual chart type also provide substantial evidence of the model’s effectiveness across various chart types.
+However, we observed that bar charts yield slightly lower metrics than other chart types.
+This discrepancy may be attributed to the inherent characteristics of bar charts, which utilize separated bars to encode data, resulting in less distinctiveness compared to the continuous line marks used in line charts and the filled area marks used in area charts.
+To enhance the accuracy of bar charts, factors such as bar widths may need to be considered during fine-tuning process.
+The comparisons between fine-tuned CLIP and zero-shot CLIP models confirm that fine-tuning is crucial for enhancing model performance in our specific domain.
+While zero-shot OpenCLIP slightly outperforms zero-shot CLIP, fine-tuned CLIP significantly surpasses fine-tuned OpenCLIP.
+Zero-shot OpenCLIP outperforms zero-short CLIP is probably due to more extensive pre-training data that include more diverse chart images.
+However, after fine-tuning, CLIP’s architecture process is better suited for our specific requirements.
+Its superior results validate our choice of CLIP as the base model.
+Figure 7: Ratings of the visual similarity between pairs of the matched sketches and charts across our fine-tuned MLLM and three pre-trained models based on a 5-point Likert scale.
 6.1.3 Sketch-Chart Retrieval
 We validate the alignment for chart-sketch pairs through a user study.
-We selected three pre-trained models, CLIP[41], OpenCLIP[8], and UniChart[37] for comparison.
 Initially, we randomly selected 200 sketches from the chart-sketch dataset.
-Then, we tasked each of the four models with determining the closest chart in the intent space to each selected sketch, indicating the chart with the highest similarity.
-This process constructed 800 matched sketch-chart pairings for subsequent user study.
+Then, we tasked each of the four models (ours, CLIP, OpenCLIP, and UniChart) with determining the closest chart in the intent space to each selected sketch, indicating the chart with the highest similarity.
+This process constructed 800 matched sketch-chart pairings for the subsequent user study.
 We recruited 30 participants (15 women, and 15 men) with experience in building charts.
@@
 The comparison results underscore the importance of encoding patterns in our alignment model, confirm the judicious choice of CLIP-ViT as our base pre-trained model, and highlight the superiority of our fine-tuned model in binding sketch and chart modalities.
-Figure 6: Ratings of the visual similarity between pairs of the matched sketches and charts across our fine-tuned LVLM and three pre-trained models based on a 5-point Likert scale.
 6.2 Case Study 1: Financial Data Exploration
@@
 Mike first loads the DJI.
-data into VisLTR and selects ‘Close’ as the displayed variable in the Main View.
-VisLTR initially generates 64,261 visualization references, among which 13,823 visualizations are pruned for exploration.
+data into VisTR and selects ‘Close’ as the displayed variable in the Main View.
+VisTR initially generates 64,261 visualization references, among which 13,823 visualizations are pruned for exploration.
 The average retrieval time for querying a visualization reference stored in Chroma is within 0.6s.
@@
   • Table Summarization.
-Mike first enters in the dialogue box “Describe the data.”.
-The response from VisLTR includes two parts, one is the basic table description “The table shows the market performance of the Dow Jones Index from January 2, 2020 to December 29, 2023, including the date, opening index, intraday high index, intraday low index, and closing index.”, and the other is the overall trend description “The data suffers a sharp drop and then a gradual recovery.
+Mike first enters the dialogue box “Describe the data.”.
+The response from VisTR includes two parts, one is the basic table description “The table shows the market performance of the Dow Jones Index from January 2, 2020, to December 29, 2023, including the date, opening index, intraday high index, intraday low index, and closing index.”, and the other is the overall trend description “The data suffers a sharp drop and then a gradual recovery.
 There are lots of fluctuations during the overall upward trend.”.
@@
 Given this description, Mike is interested in the “sharp drop” pattern, so he tries to locate the pattern by entering the text “Please give me more details about the mentioned ‘sharp drop’ pattern.”.
-Based on the context, VisLTR can easily locate the “sharp drop” pattern appears, and the matched interval is highlighted in Main View.
+Based on the context, VisTR can easily locate the “sharp drop” pattern that appears, and the matched interval is highlighted in Main View.
 At the same time, a text description “In March 2020, the Dow Jones Index experienced a ‘sharp drop’ pattern.
@@
 Mike then freely sketches a ‘w’ pattern [Uncaptioned image] in the empty canvas and enters “Are there patterns in the data similar to my hand-drawn sketch?”.
-VisLTR retrievals for the visualization references similar to the input hand-drawn sketch from the vector database, and returns the three most similar matches in Main View.
-To facilitate user understanding, VisLTR also returns a text description of the matched results: “The three intervals with similar patterns are from 8-Feb-21 to 16-Mar-21, 18-Feb-21 to 18-Mar-21, and 15-Dec-20 to 11-Jan-21.
+VisTR retrievals for the visualization references similar to the input hand-drawn sketch from the vector database, and returns the three most similar matches in Main View.
+To facilitate user understanding, VisTR also returns a text description of the matched results: “The three intervals with similar patterns are from 8-Feb-21 to 16-Mar-21, 18-Feb-21 to 18-Mar-21, and 15-Dec-20 to 11-Jan-21.
 This pattern is recognized as a double-bottom pattern.
@@
 In this way, Mike gets an overall understanding of the tabular data, and also typical trends and patterns within the data.
-The multi-modal interactions, together with visual interactions like selecting and zooming in the visual interface, enable Mike to easily understand the DJI.
-data and identify interesting data change patterns.
-Figure 7: Case study 1: data exploration for DJI. form 2020 to 2023. Mike gets a complete understanding of the data via table summarization, question & answering, and visual query enabled by VisLTR.
-Figure 8: Case study 2: multivariate analysis for the uploaded air pollutant dataset. With VisLTR, Emily identifies typical data change patterns and relationships between different air pollutants.
+The multimodal interactions, together with visual interactions like selecting and zooming in the visual interface, enable Mike to easily understand the DJI.
+data and recognize interesting data change patterns.
+Figure 8: Case study 1: data exploration for DJI. form 2020 to 2023. Mike gets a complete understanding of the data via table summarization, question & answering, and visual query enabled by VisTR.
+Figure 9: Case study 2: multivariate analysis for the uploaded air pollutant dataset. With VisTR, Emily recognizes typical data change patterns and relationships between different air pollutants.
 6.3 Case Study 2: Air Pollutant Analysis
@@
 A student, Emily, is seeking the correlations between different air pollutants.
-The used air pollutant dataset includes the daily air quality metrics from January 1, 2017 to December 31, 2018, which is collected by the Beijing US Embassy [^1] https://aqicn.org/city/beijing/us-embassy/cn/.
+The used air pollutant dataset includes the daily air quality metrics from January 1, 2017, to December 31, 2018, which is collected by the Beijing US Embassy [^1] https://aqicn.org/city/beijing/us-embassy/cn/.
 The data includes variables of Date, PM25, O3, temperature, and relative humidity, which provides insights into the fluctuations in air quality and correlations between air pollutants.
-Emily first loads the pollutant data into VisLTR, which generates a multi-line chart with each line representing a variable.
-In the backend, VisLTR generates a total of 102,787 visualization references, of which 40,243 are stored in Chroma after pruning.
-Figure 8 illustrates Emily’s analysis process using VisLTR to explore the multivariate data.
+Emily first loads the pollutant data into VisTR, which generates a multi-line chart with each line representing a variable.
+In the backend, VisTR generates a total of 102,787 visualization references, of which 40,243 are stored in Chroma after pruning.
+Figure 9 illustrates Emily’s analysis process using VisTR to explore the multivariate data.
   • Free Exploration.
-To obtain the typical data change patterns of pollutants, Emily selects four pollutant indicators, PM25, O3, temperature, and relative humidity, and obtains their typical change patterns in the Pattern View.
-She notices that all the four indicators exhibit obvious daily fluctuations.
+To obtain the typical data change patterns of pollutants, Emily selects four pollutant indicators, PM25, O3, temperature, and relative humidity, and obtains the change patterns in the Pattern View.
+She notices that all four indicators exhibit obvious daily fluctuations.
 Specifically, PM25 exhibits a ‘zigzag pattern with varying amplitudes.’, indicative of its dynamic nature and daily variability.
@@
 […]
-In this manner, Emily gains a comprehensive insight into the multivariate air pollutant data, identifying both the typical data change patterns and correlations among various air pollutant indicators.
+In this manner, Emily gains a comprehensive insight into the multivariate air pollutant data, recognizing both the typical data change patterns and correlations among various air pollutant indicators.
 7 Discussion
-The case studies have demonstrated VisLTR’s capability in identifying and explaining data change patterns within tabular data.
-By integrating visualization-in-the-loop, VisLTR not only complements SQL-based LLMs for table reasoning but also aligns with the exploratory data analysis (EDA) process.
-EDA often involves summarizing the main characteristics of a dataset using visualization, enabling users to understand the data’s structure, identify patterns, detect anomalies, and formulate hypotheses for further investigation.
+The case studies have demonstrated VisTR’s capability in recognizing and explaining data change patterns within tabular data.
+By leveraging visualizations as representation, VisTR successfully addresses the limitations of existing LLM-based table reasoning methods in pattern recognition and visual-based pattern exploration.
+Moreover, the process also aligns with the exploratory data analysis (EDA) process.
+EDA often involves summarizing the main characteristics of a dataset using visualization, enabling users to understand the data’s structure, recognize patterns, detect anomalies, and formulate hypotheses for further investigation.
 In this sense, the proposed framework not only streamlines the process of EDA but also expands its horizons, providing users with novel approaches to uncover hidden patterns within datasets.
-Generalizability. Nevertheless, to fully realize the potential of the proposed framework, we need to enhance the generalizability of VisLTR from the following perspectives.
+Generalizability. Nevertheless, to fully realize the potential of the proposed framework, we need to enhance the generalizability of VisTR from the following perspectives.
   • More table types.
-VisLTR is currently instantiated on time-series tables, which contain data change patterns that cannot be processed by existing SQL-based LLMs for table reasoning.
-VisLTR has the potential to be extended to more general table types by employing alternative data transformation methods within the visualization referencing module.
-For instance, VisLTR can be adapted to recognize and analyze the distribution patterns in a student grade table.
-In this scenario, a data aggregation method could be employed to identify grade distribution patterns, such as the normal distribution.
+VisTR is currently instantiated on time-series tables, which contain data change patterns that cannot be processed by existing SQL-based LLMs for table reasoning.
+VisTR has the potential to be extended to more general table types by employing alternative data transformation methods within the visualization referencing module.
+For instance, VisTR can be adapted to recognize and analyze the distribution patterns in a student grade table.
+In this scenario, a data aggregation method could be employed to recognize grade distribution patterns, such as the normal distribution.
 Supporting a broader range of table types is feasible, leveraging methods such as subset embedding[62] and learning-based approaches[54].
   • General user queries.
-VisLTR focuses on tasks related to reasoning about data change patterns in tables.
-Therefore, we primarily provide text-based user queries related to data change patterns, such as “what type of change pattern exists in the data?” However, this narrow focus also limits our ability to handle other general user queries, including those related to causal reasoning, such as “why did Apple stock rise in March?” When posed with such questions, VisLTR may encounter the ‘hallucination’ issue, leading to incorrect queries and responses.
+VisTR focuses on tasks related to reasoning about data change patterns in tables.
+Therefore, we primarily provide text-based user queries related to data change patterns, such as “what type of change pattern exists in the data?” However, this narrow focus also limits our ability to handle other general user queries, including those related to causal reasoning, such as “why did Apple stock rise in March?” When posed with such questions, VisTR may encounter the ‘hallucination’ issue, leading to incorrect queries and responses.
 To achieve a comprehensive exploration and reasoning of a data table, we require more general user queries, which may necessitate the incorporation of external knowledge sources.
@@
 Users’ intuition about sketching targets often emerges from the exploration process.
-As a reason, VisLTR offers insights into common patterns within the uploaded table.
+As a reason, VisTR offers insights into common patterns within the uploaded table.
 However, expressing complex and less common patterns solely through hand-drawn sketches can be challenging for users[32].
@@
 Summarizing multi-attribute temporal patterns and introducing a more effective time-series segmentation method have shown promise in addressing this issue[51].
-By implementing these methods, we can further enhance the capabilities of VisLTR to support sketches with enhanced intents.
+By implementing these methods, we can further enhance the capabilities of VisTR to support sketches with enhanced intents.
 Scaling up the dataset.
 The successful alignment of our model heavily relies on acquiring high-quality training data.
-Scaling up the dataset to enhance the generalizability of VisLTR presents significant challenges for the data augmentation and user labeling strategies currently employed.
+Scaling up the dataset to enhance the generalizability of VisTR presents significant challenges for the data augmentation and user labeling strategies currently employed.
 Collecting user-labeled sketches is time-consuming, and incorporating intents into sketches would further increase complexity.
@@
 Besides table reasoning, the proposed framework holds potential for other application scenarios, such as safeguarding sensitive data that requires a balance between data privacy and utility.
-By transforming raw tabular data into visualizations, VisLTR shields the secure information contained within tables from users, while still allowing for examination and querying of the data patterns.
-Moreover, VisLTR could also find applications in fraud detection, as visualizations generated from transactional data could help identify suspicious patterns indicative of fraudulent activities.
+By transforming raw tabular data into visualizations, VisTR shields the secure information contained within tables from users, while still allowing for examination and querying of the data patterns.
+Moreover, VisTR could also find applications in fraud detection, as visualizations generated from transactional data could help recognize suspicious patterns indicative of fraudulent activities.
 8 Conclusion
-This work introduces VisLTR system built upon a novel visualization-in-the-loop framework for table reasoning, serving as a complement to existing SQL-based table reasoning research.
-By utilizing visualizations as a proxy, the framework offers several advantages, including the ability to identify data change patterns, mitigate LLM drift, and support multi-modal interactions.
-The implementation of VisLTR requires an effective multi-modal visualization alignment model that binds chart, text, and sketch modalities together in a joint embedding space.
-This is accomplished by fine-tuning an LVLM with a new dataset of chart-text and chart-sketch pairings.
-Together with strategies employed for visualization referencing, indexing, and retrieval, we have created an interactive system that facilitates multi-modal interactions with tabular data.
-The feasibility of VisLTR is demonstrated through quantitative evaluations of the visualization alignment model and practical case studies.
-The proposed visualization-in-the-loop framework offers a new research direction for table reasoning.
-We aim to extend the support to a wider range of table types, accommodate more generalized user queries, and explore additional application scenarios in the future.
+This work introduces VisTR system, built upon a novel framework for table reasoning, expanding upon existing LLM-based table reasoning research.
+By leveraging visualizations as representations, the framework offers several advantages, including the ability to recognize data change patterns, mitigate LLM drift, and support visual-based exploration.
+The implementation of VisTR requires an effective multimodal visualization alignment model that binds chart, text, and sketch modalities together in a joint embedding space.
+This is accomplished by fine-tuning an MLLM with a new dataset of chart-text and chart-sketch pairings.
+Combined with strategies employed for visualization referencing, indexing, and pruning, we have developed an interactive system that facilitates multimodal interactions with time-series tables.
+The feasibility of VisTR is demonstrated through quantitative evaluations of the visualization alignment model and practical case studies on time-series data tables.
+This proposed framework offers a new research direction for table reasoning and enhances the fuzzy exploration process of existing LLM-based table reasoning methods.
+We aim to extend its support to a wider range of table types, accommodate more generalized user queries, and explore additional application scenarios in the future.
 Acknowledgements.

### v3

@@
-VisTR: Visualizations as Representations for Time-series Table Reasoning
+mO Yuyu[#1]
 Abstract
-Table reasoning involves transforming natural language questions into corresponding answers based on the provided data table.
-Recent research exploits large language models (LLMs) to facilitate table reasoning, which however struggle with pattern recognition and lack support for visual-based pattern exploration.To address these limitations, we propose VisTR, a framework that leverages visualizations as representations to facilitate data pattern recognition and support cross-modal exploration.
-We describe VisTR as a process consisting of four major modules: 1) visualization alignment that utilizes multimodal LLMs to align visualizations across various modalities, including chart, text, and sketch; 2) visualization referencing that decomposes a table into multifaceted visualization references that comprehensively represent the table; 3) visualization pruning that incorporates data and retrieval pruning to excise visualization references with poor information and enhance retrieval efficiency; and 4) visualization interaction that offers an interactive visual interface with multimodal interactions for user-friendly table reasoning.
-Quantitative evaluation with existing multimodal LLMs demonstrates the effectiveness of the alignment model in cross-modal visualization pairings.
-We further illustrate the applicability of the proposed framework in various time-series table reasoning and exploration tasks.
-keywords:
-Table reasoning, visualization representation, multimodal LLMs
-[…]
-VisTR leverages visualizations as the representations between a table and user intentions.
-The proposed framework advances existing LLM-based table reasoning methods by supporting data change pattern recognition (Q1), improving pattern recognition for long tables (Q2), and enabling visual-based pattern exploration (Q3).
-1 Introduction
-Table reasoning is a crucial natural language processing task that involves automatically extracting and inferring information from tables to answer specific questions[12, 70].
-This task supports a wide range of applications such as fact verification[6] and question answering[26], where each question-table pair constitutes an instance for reasoning[69].
-Given the superior commonsense and logical reasoning capabilities of large language models (LLMs)[63], utilizing them for table reasoning has emerged as the predominant approach[5, 12].
-Existing methods can be broadly categorized into generic table reasoning, where pre-trained models or LLMs directly generate answers, often enhanced by techniques such as designing prompts[5] to guide iterative reasoning or employing flexible tabular formats[57, 69]; and program-aided table reasoning that utilizes executable query languages (e.g., SQL, SPARQL, or Python) as the bridge between user queries and answers derived from the table[65, 9, 19].
-Despite their advancements, there exist several limitations.
-First, these methods struggle with pattern recognition, which involves the automatic identification and extraction of meaningful patterns within data, such as trends in time-series data[35, 8].
-For instance, Q1 in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning illustrates a scenario where users seek to examine the price trend of Apple during March.
-Existing LLM-based methods cannot correctly capture and respond to the data change patterns within the corresponding time period.
-The challenge becomes more pronounced when dealing with large-size tables, as LLMs are known to encounter ‘drift’ issues, where models lose information in the middle of long contexts[4].
-This loss of context can result in errors during the reasoning process, making it difficult to extract patterns accurately, as illustrated in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q2).
-Last, the methods lack support for visual-based pattern exploration, making it challenging for users to iteratively engage with data tables, particulalry for lay users.
-Rather than expressing their intentions precisely in natural language, laymen would prefer to use multimodal inputs like hand-drawn sketches or charts[41].
-An example is shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q3), where a user may not know the term “two-valley” but can easily recognize this visual pattern from the collected bar chart.
-However, existing LLM-based methods do not support such visual-based exploration, which are increasingly in demand.
-To overcome the limitations of existing LLM-based table reasoning methods, we introduce a novel framework that incorporates visualization as representation into multimodal LLMs (MLLMs).
-Our key idea is to transform data tables into a set of meaningful and insightful visualizations, termed visualization references, which serve as a concise data representation to capture data insights and align with human intentions.
-This framework offers several advantages that address the shortcomings of existing LLM-based table reasoning methods, effectively enhancing the table exploration process.
-First, visualization references can help recognize the short- and long-term data change patterns.
-For instance, when recognizing price trend in the given data table (Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q1)), the corresponding visualization reference could capture the “two-peak” price trend of Apple during March.
-Second, regardless of the length of data facets, the visualization references can capture and retain the data change patterns in fixed-size images.
-Adopting fixed-size visualization references as the representations of data facets can effectively circumvent the ‘drift’ issues encountered during reasoning.
-For Q2 in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning, the stock price change over a year can be represented as a single visualization, where the price bottom can be easily located in this reference.
-Moreover, leveraging visualizations as a representation offers new modalities for interacting with the tabular data, aligning more intuitively with human intentions[32].
-Combined with multimodal input, including text and chart (as shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(Q3)), it becomes easier to explore data tables than directly observing the raw table.
-We illustrate the effectiveness of VisTR on time-series data tables, which often feature large table sizes and complex data change patterns.
-In this context, the proposed framework also presents several challenges to be addressed.
-First, time-series data often contain a variety of long- and short-term patterns, which should be comprehensively captured by the visualization references.
-To ensure the visualization references are comprehensive enough to accurately represent a data table, thereby meeting the diverse analytical needs, poses the first challenge (C1).
-Second, to effectively align textual descriptions, visualization images, and hand-drawn sketches to bridge user intentions with visualization references, meanwhile enabling multimodal interactions presents a challenge (C2).
-Existing models for vision-language tasks such as chart question answering (CQA)[24, 68] are not specifically designed to meet this requirement, thus being less effective in identifying data change patterns.
-Last, to comprehensively represent a table, a large number of visualization references will be generated.
-Effectively indexing the visualization references and supporting efficient visualization retrieval for timely user response present a significant challenge (C3).
-We design a prototype system that encompasses four main modules aimed at addressing the challenges C1--C3.
-First, the visualization referencing module decomposes an input table into data facets and processes the facets into corresponding visualization references that comprehensively capture the essence of the table.
-To ensure a comprehensive representation of data patterns in tabular data, we adopt a multifaceted process to get meaningful facets and then transform each facet into a fixed-size visualization reference (addressing C1).
-Second, in the visualization alignment module, we fine-tune a MLLM to connect visualizations of various modalities including charts, text, and sketches (addressing C2).
-Specifically, we construct a new cross-modal visualization dataset using data augmentation for chart-text pairing and user labeling for chart-sketch pairing, taking into account user intentions.
-We further leverage the CLIP model[48] enriched with two-level cross-entropy loss and bidirectional hinge-based triplet ranking loss to enhance the alignment.
-Next, the visualization pruning module leverages visualization pruning to filter out less informative visualization references, and visualization indexing to enhance the efficiency of the retrieval process for the embeddings of visualization references stored in a vector database (addressing C3).
-Finally, we design a visualization interface that enables users’ exploration of time-series data tables through multimodal interactions.
-Here, we incorporate a “decomposing-execution-filling” strategy to integrate the output generated by VisTR with the input table.
-Specifically, the user input is initially decomposed into visualization retrieval tasks, which are executed in the vector database, and their corresponding text descriptions are extracted by the fine-tuned MLLM.
-Subsequently, these descriptions are filled back into the input table to generate textual responses for the user.
-Quantitative evaluations demonstrate the effectiveness of the fine-tuned MLLM in aligning different modalities.
-Additionally, case studies illustrate how the proposed framework addresses the limitations of existing LLM-based table reasoning methods through various time-series table reasoning and visual-based exploration tasks.
-Dataset, fine-tuned model, and evaluation results are released at https://osf.io/yun9d/.
+Time-series table reasoning interprets temporal patterns and relationships in data to answer user queries.
+Despite recent advancements leveraging large language models (LLMs), existing methods often struggle with pattern recognition, context drift in long time-series data, and the lack of visual-based reasoning capabilities.
+To address these challenges, we propose VisTR, a framework that places visualizations at the core of the reasoning process.
+Specifically, VisTR leverages visualizations as representations to bridge raw time-series data and human cognitive processes.
+By transforming tables into fixed-size visualization references, it captures key trends, anomalies, and temporal relationships, facilitating intuitive and interpretable reasoning.
+These visualizations are aligned with user input, i.e., charts, text, and sketches, through a fine-tuned multimodal LLM, ensuring robust cross-modal alignment.
+To handle large-scale data, VisTR integrates pruning and indexing mechanisms for scalable and efficient retrieval.
+Finally, an interactive visualization interface supports seamless multimodal exploration, enabling users to interact with data through both textual and visual modalities.
+Quantitative evaluations demonstrate the effectiveness of VisTR in aligning multimodal inputs and improving reasoning accuracy.
+Case studies further illustrate its applicability to various time-series reasoning and exploration tasks.
+Index Terms:
+Table reasoning, visualization representation, multimodal LLMs.
+publicationid: pubid: $0$–$00$ 2021 IEEE
+I Introduction
+Figure 1: VisTR leverages visualizations as the representations to bridge the gap between time-series data and user queries.
+VisTR enhances existing LLM-based table reasoning methods by enabling robust data change pattern recognition (Q1), improving pattern recognition for long-term series (Q2), and facilitating visual-based pattern exploration (Q3).
+Time-series tabular data are ubiquitous, forming the backbone of decision-making processes across domains such as finance, healthcare, and logistics[1, 2].
+These datasets record sequential information over time, enabling the discovery of temporal trends and interesting relationships that drive actionable insights[3, 4].
+Time-series table reasoning, a key task in this context, focuses on interpreting these patterns to answer user queries, such as “What is the price trend of Apple during March?” (see Fig.
+1-Q1).
+Intuitively, effectively addressing this task requires a comprehensive understanding of not only the raw tabular data but also the temporal trends, interrelationships among data points, and alignment with the specific context of user queries [5].
+Recently, large language models (LLMs) have demonstrated promising capabilities in extending their applications to table reasoning tasks [6, 7, 8].
+These methods can be broadly categorized into: (1) generic table reasoning, where pre-trained models or LLMs directly generate answers, often enhanced by techniques such as prompt design [9, 10], iterative reasoning, or flexible tabular formats [11, 12]; and (2) program-aided table reasoning which employs executable query languages (e.g., SQL, SPARQL, or Python) to bridge user queries with answers derived from tabular data [13, 14, 15].
+Limitations of Existing Solutions. Despite these advances, existing solutions exhibit limitations, as illustrated in Fig. 1.
+(1) Limited Pattern Recognition.
+LLMs often struggle to identify and interpret temporal data patterns, such as trends, peaks, or valleys [1, 16].
+For example, in Q1 of Fig.
+1, an LLM-based approach misinterprets Apple’s stock prices as exhibiting a “declining trend” rather than recognizing the correct two-peak trend.
+(2) Challenges with Long Time-series Data.
+LLMs often lose focus and context when reasoning over long time-series data, leading to errors caused by “LLM drift” [17].
+In Q2 of Fig.
+1, an LLM-based method erroneously identifies Google’s stock price bottom as occurring “at the beginning” rather than pinpointing the correct date.
+(3) Inability to Perform Visual-based Reasoning.
+Existing LLMs cannot effectively reason over external visual references, limiting their ability to align time-series data with visual patterns [18].
+For example, in Q3 of Fig.
+1, an LLM-based method fails to determine whether Tesla’s stock price exhibits a similar pattern to a user-provided chart, demonstrating the gap in visual-based reasoning capabilities.
+Rethinking: How Can Visualization Help?
+Visualizations, such as line and bar charts, offer an intuitive means of representing time-series data by highlighting trends, outliers, and temporal relationships [19, 20].
+They act as a cognitive bridge, enabling humans to discern patterns that might otherwise remain obscured in raw time-series data.
+Despite their potential, existing table reasoning methods largely overlook the use of visualizations as proxies for reasoning.
+Instead, they rely heavily on text-based reasoning, which struggles to emulate human-like cognitive processes and often fails in tasks requiring complex pattern recognition and context alignment with user intentions.
+Our Proposal.
+Inspired by the ability of visualizations to simplify complex patterns and align with human cognitive processes, we propose a new framework that integrates visualizations as representations into multimodal large language models (MLLMs).
+Our goal is to achieve the best of both worlds by combining the interpretability and intuitive pattern recognition offered by visualizations with the reasoning and generative capabilities of LLMs, thereby enhancing both the accuracy and usability of time-series table reasoning systems.
+The key idea of our framework is to transform time-series data into a set of meaningful and insightful visualizations, termed visualization references.
+These references act as concise representations of the data, capturing essential insights such as trends, anomalies, and temporal relationships, while aligning closely with user intentions for reasoning tasks.
+Framework Overview.
+At its core, the framework transforms input tables into fixed-size visualization references that capture key trends, anomalies, and temporal relationships, enabling a comprehensive and interpretable representation of both long- and short-term data patterns.
+To achieve this, the framework begins by decomposing tables into smaller, meaningful data facets, each representing a subset of the table’s temporal or structural characteristics.
+These facets are then converted into visualizations, ensuring a rich and diverse representation of the data.
+Second, to align user intentions with these visualization references, the framework incorporates a fine-tuned MLLM that constructs a unified embedding space for multimodal inputs, including charts, text, and hand-drawn sketches.
+A novel cross-modal dataset is developed to enhance alignment precision, combining augmented chart-text pairs with user-labeled chart-sketch pairs.
+By extending the CLIP model [21] with advanced loss functions, e.g., a two-level cross-entropy loss and a bidirectional triplet ranking loss, the framework ensures accurate and consistent alignment across modalities, bridging the gap between user inputs and relevant visualizations.
+Third, given the large-scale visualization references generated from extensive datasets, the framework incorporates a pruning and indexing mechanism to enhance scalability and retrieval efficiency.
+Less informative visualizations are filtered out, while the remaining embeddings are indexed in a vector database, enabling fast and accurate retrieval during user interactions.
+Finally, the framework integrates these components into an intuitive visualization interface supporting seamless multimodal interaction.
+Using a “decomposing-execution-filling” strategy, user queries are broken into retrieval tasks, linked to textual descriptions generated by the fine-tuned MLLM, and reintegrated into the table for final responses.
+This interface enables interactive exploration through charts, text, and sketches.
+Our framework offers several key advantages over existing LLM-based table reasoning methods.
+(1) Enhanced Pattern Recognition Across Time Scales.
+Visualization references effectively capture both short- and long-term data change patterns, making it easier to identify nuanced temporal trends.
+For example, in Fig.
+1(Q1), the “two-peak” price trend of Apple during March is clearly represented, overcoming the limitations of purely text-based approaches.
+(2) Mitigating Context Drift in Long Time-series.
+Fixed-size visualization references condense data patterns, mitigating context drift issues in long time-series reasoning.
+For example, in Fig.
+1(Q2), a year-long stock price change is condensed into a single visualization, making it easy to pinpoint the price bottom without losing contextual accuracy.
+(3) Intuitive Multimodal Interaction.
+Integrating visualizations as representations enables natural, multimodal data exploration by aligning with human cognitive processes.
+For example, Fig.
+1(Q3) shows how combining textual descriptions with charts improves query interpretation and reduces reliance on raw time-series data [22].
 In summary, we make the following contributions:
-(1) Framework. We introduce a framework that leverages visualizations as the representations to enhance table reasoning, to facilitate data change pattern recognition and visual-based exploration.
-(2) Multimodal Visualization Alignment.
-We design an effective MLLM that learns a joint embedding space across three visualization modalities - charts, text, and hand-drawn sketches.
-The model aligns multimodal user intentions with visualization references, allowing users to explore and analyze tables through various interactions.
-(3) VisTR System.
-We develop a prototype system that supports the exploration and reasoning of large-scale and multiple time-series tables through intuitive user interactions.
-Various case studies and quantitative results prove the effectiveness and applicability of VisTR.
-2 Related Work
-Table Reasoning.
+  • Framework.
+We propose VisTR that leverages visualizations as representations to enhance time-series table reasoning.
+This new framework enables robust pattern recognition and intuitive exploration for time-series data.
+  • Multimodal Visualization Alignment.
+We finetune an effective MLLM that learns a joint embedding space across three modalities, i.e., charts, text, and hand-drawn sketches.
+This alignment bridges user intentions with visualization references, enabling users to explore time-series tables through multimodal interactions.
+  • Prototype Construction. We implement a prototype system that facilitates reasoning and exploration of large-scale time-series tables.
+  • Evaluation.
+We conduct extensive evaluations, including quantitative experiments and case studies, to demonstrate the effectiveness and usability of VisTR.
+All datasets, fine-tuned models, and evaluation results are available at https://osf.io/yun9d/.
+II Related Work
+II-A Table Reasoning
 Table reasoning integrated with natural language (NL) interfaces enables lay users to explore tabular data effortlessly, supporting complex tasks such as fact verification[23] and question answering[24].
@@
 Figure 2: An overview of the proposed framework, which mainly contains four main modules of visualization referencing, visualization pruning, visualization alignment, and visualization interaction.
-Multimodal Visualization Alignment.
+II-B Multimodal Visualization Alignment
 Given that visualizations are inherently stored as images, we opt to leverage emerging MLLMs to embed visualization images and facilitate multimodal user interactions.
 In recent years, significant advancements have been made in utilizing vision-language models for chart-based tasks, such as CQA [35, 36, 37], chart retrieval [38] and generation [39, 40].
-Though CQA models can provide answers for user utterances, these models only support text-chart interactions, lacking the capability for cross-modal interactions involving multiple modalities such as text, chart, and sketch preferred by laymen.
+Though CQA models can provide answers for user utterances, these models only support text-chart interactions, lacking the capability for cross-modal interactions involving multiple modalities such as text, chart, and sketch preferred by laymen [41].
 Furthermore, these models are not specifically designed for effectively recognizing data change patterns (see Sect.
-6.1), a gap that this work seeks to fill in the context of table reasoning.
+VI-A), a gap that this work seeks to fill in the context of table reasoning.
 To this end, we choose to construct a novel vision-language model with an innovative visualization alignment strategy.
@@
 To address this gap, we utilize data augmentation and user labeling strategies to supplement existing datasets in training the multimodal visualization alignment model.
-Visualization Storage, Indexing, and Retrieval.
+II-C Visualization Storage, Indexing, and Retrieval
 To comprehensively represent tabular data, our approach will generate a large number of visualization references and encode them into vector representations, thus raising new challenges for efficiently storing, indexing, and retrieving them.
@@
 The combination of pruning with indexing allows for quick retrieval and handling timely user queries.
-3 The Visualization Enhanced Framework
-The visualization enhanced framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Figure 1.
+III The Visualization Enhanced Framework
+The visualization enhanced framework consists of four modules: visualization alignment, visualization referencing, visualization pruning, and visualization interaction, as shown in Fig. 2.
 Visualization Alignment.
@@
 A detailed description of the module is provided in Sect.
-4.
+IV.
 Visualization Referencing.
@@
 The description of this module is detailed in Sect.
-5.1.
+V-A.
 Visualization Pruning.
@@
 This module is detailed in Sect.
-5.2.
+V-B.
 Visualization Interaction.
-In the visualization interaction module, we introduce a “decomposing-execution-filling” strategy to enable the connection of multimodal user queries with the stored embeddings of visualization references.
+For visualization interaction, we introduce a “decomposing-execution-filling” strategy to enable the connection of multimodal user queries with the stored embeddings of visualization references.
 Specifically, the user query is first decomposed into a visualization retrieval task.
-Second, the retrieval task is executed in Chrome to get a retrieved visualization reference.
+Second, the retrieval task is executed in Chroma to get a retrieved visualization reference.
 Afterward, with the help of the fine-tuned MLLM, the corresponding trend word of the retrieved visualization reference is generated and then filled into the textual response to the user.
 An interactive visual interface with multiple views and rich interactions is also developed to support the iterative table reasoning process, transitioning from abstract to detailed exploration, as detailed in Sect.
-5.3.
-4 Multimodal Visualization Alignment
+V-C.
+IV Multimodal Visualization Alignment
 We set out to establish a MLLM that aligns visualization modalities of chart, sketch, and text in a single joint embedding space.
 This section presents the data preparation, including data augmentation and user labeling strategies, to address the issue of lacking training dataset (Sect.
-4.1), followed by modality binding for establishing a joint embedding space that binds chart, sketch, and text modalities together (Sect.
-4.2).
+IV-A), followed by modality binding for establishing a joint embedding space that binds chart, sketch, and text modalities together (Sect.
+IV-B).
 In the end, we present the implementation details (Sect.
-4.3).
+IV-C).
 Figure 3: An overview of visualization alignment module, which contains data preparation for chart-text and chart-sketch pairings, and achieves the alignment for chart, sketch, and text modalities.
-4.1 Data Preparation
+IV-A Data Preparation
 To train the multimodal visualization alignment model, we need to first prepare a dataset containing user intentions, which is not covered by current open-sourced datasets.
 To achieve the alignment of sketches, charts, and text, we employ two strategies to construct the training data: (1) data augmentation for chart-text pairing, and (2) user labeling for chart-sketch pairing.
-Figure 2 uses an example to illustrate the two data preparation strategies.
+Fig.
+3 uses an example to illustrate the two data preparation strategies.
 (1) Data augmentation for chart-text pairing.
@@
 Figure 4: The alignment pipeline for three modalities. Each modality is processed by a separate encoder, and the chart modality is set as the supervisor, guiding the alignment process.
-4.2 Modality Binding
+IV-B Modality Binding
 We aim to bind all text $\mathbb{T}$, chart $\mathbb{C}$ and sketch $\mathbb{S}$ modalities together in a joint embedding space.
 Existing CLIP-based works predominantly focus on aligning pairs of modalities, while our objective is to accomplish the alignment of three modalities in the absence of the pairings between text and sketch modalities.
-As shown in Figure 3, we employ a two-stage strategy: first, we employ modality-specific encoders to transform each modality into embeddings of uniform length[59].
+As shown in Fig.
+4, we employ a two-stage strategy: first, we employ modality-specific encoders to transform each modality into embeddings of uniform length[63].
 All three modalities are processed using the Transformer architecture, leveraging its self-attention mechanism to effectively capture complex dependencies within each modality.
@@
 The formulation for the two cross-entropy functions is specified as:
-$H_{1}(C,T)=-\sum(p_{i}\times\log(k_{i})),$
-$H_{2}(C,S)=-\sum(p_{i}\times\log(q_{i})),$
+$H_{1}(C,T)=-\sum(p_{i}\times\log(k_{i})),$ (1)
+$H_{2}(C,S)=-\sum(p_{i}\times\log(q_{i})),$ (2)
 where $H_{1}$ calculates the difference between charts $C$ and text $T$, and $(C,T)$ is the matched positive chart-text pair.
@@
 Take chart and sketch modalities as an example, the loss is as follows:
-$L_{2}(C,S)=[\alpha+H_{2}(C^{\prime},S)-H_{2}(C,S)]_{+}+[\alpha+H_{2}(C,S^{\prime})-H_{2}(C,S)]_{+},$
+$L_{2}(C,S)=[\alpha+H_{2}(C^{\prime},S)-H_{2}(C,S)]_{+}$
+$+[\alpha+H_{2}(C,S^{\prime})-H_{2}(C,S)]_{+},$ (3)
 where $\alpha$ denotes the margin parameter, and $[x]_{+}\equiv\max(x,0)$.
@@
 Set $L$ as the final loss of three modalities in the embedding space, that is, the sum of the two ranking losses: $L(T,C,S)=L_{1}(C,T)+L_{2}(C,S).$ By minimizing both the loss between chart and text modalities, and the loss between chart and sketch modalities, we achieve the alignment between the similar embeddings of the three modalities in a joint embedding space.
-4.3 Implementation Details
+IV-C Implementation Details
 We took a pre-trained CLIP text-image model and fine-tuned it on the selected chart-text and chart-sketch pairings.
@@
 This behavior is also validated by IMAGEBIND[45] so we reference to its statement that $(\mathbb{C},\mathbb{S})$ and $(\mathbb{C},\mathbb{T})$ are naturally aligned, and the text $\mathbb{T}$ and sketch modalities $\mathbb{S}$ constitute an emergency alignment.
-5 VisTR System
+V VisTR System
 With the MLLM for multimodal visualization alignment, we build VisTR system that allows users to intuitively explore, comprehend, and interact with tabular data using visualizations as the representation.
 This section details the implementation of visualization referencing (Sect.
-5.1), visualization pruning (Sect.
-5.2), and visualization interaction (Sect.
-5.3) in VisTR system.
-5.1 Visualization Referencing
+V-A), visualization pruning (Sect.
+V-B), and visualization interaction (Sect.
+V-C) in VisTR system.
+V-A Visualization Referencing
 This module is designed to ensure that the generated visualization references can effectively capture comprehensive data patterns inherent in the raw tabular data.
@@
 Collectively, for all variables in the input table $D$, the visualization references are denoted as $C=\bigcup_{i=1}^{n}C(d_{i})=\{c_{i,j}|d_{i}\in D,r_{j}(d_{i})\in R(d_{i})\}$.
-5.2 Visualization Pruning and Indexing
+V-B Visualization Pruning and Indexing
 For a given table, we will generate a vast collection of visualization references, each of which is subsequently mapped into vector representations (i.e., embeddings).
 The large number of visualization references poses significant challenges to efficient storage and retrieval.
-To address these challenges, we design a combined approach of visualization pruning and indexing, as illustrated in Figure 4, to improve VisTR’s ability to manage and retrieve visualization references.
+To address these challenges, we design a combined approach of visualization pruning and indexing, as illustrated in Fig.
+5, to improve VisTR’s ability to manage and retrieve visualization references.
 Figure 5: The pipeline of visualization referencing and visualization pruning modules. Visualization pruning and indexing are used to accelerate the storage and retrieval process.
@@
 If $d(c_{j},c_{k})$ is less than the set pruning threshold, the vector covering a shorter time range is filtered out.
-An example scenario is illustrated in Figure 4, where the ‘a peak’ change pattern exists in pink-boxed visualization references.
+An example scenario is illustrated in Fig.
+5, where the ‘a peak’ change pattern exists in pink-boxed visualization references.
 In this case, the light-pink-boxed one has less information and is consequently filtered out.
@@
 The pruning threshold acts as a predefined parameter that affects the number of stored visualization references.
-To assess its impact, we designed an experiment with four values: 0.5, 0.8, 1, and 1.5, and applied the thresholds to tables in a range of row numbers.
+To assess its impact, we designed an experiment with four values: 0.5, 0.8, 1, and 1.5, and applied the thresholds to tables of varying row counts.
 Specifically, the row numbers were considered based on the typical sizes used in existing LLM-based table reasoning methods, which typically support data tables with fewer than 50 rows.
@@
 data table for each pruning threshold.
-Figure 5 shows the relationship between the number of stored visualization references, the number of data table rows, and the pruning thresholds.
+Fig.
+6 shows the relationship between the number of stored visualization references, the number of data table rows, and the pruning thresholds.
 When the pruning threshold is set to 1, the number of stored visualization references approximated 20% of the raw reference number.
@@
 Figure 7: The interactive interface of VisTR, which contains four main views and supports multimodal interactions for table reasoning.
-5.3 Visualization Interaction
+V-C Visualization Interaction
 The Decomposing-Execution-Filling Strategy.
@@
 Here, we employ two LLMs with in-context learning using prompts to decompose user queries and complete outputs respectively.
-Below we illustrate an example using the case shown in Figure VisTR: Visualizations as Representations for Time-series Table Reasoning(A).
+Below we illustrate an example using the case shown in Fig.
+1(A).
 Given the query “What is the price trend of Apple during March?”, we first decompose the question into a retrieval task “return the visualization reference of Apple price in March”, and a cross-modal alignment task “recognize the trend in the retrieved visualization reference”.
@@
 In addition to textual responses, VisTR also incorporates a user-friendly interface to enable intuitive table reasoning and exploration.
-As shown in Figure 6, the interface includes four main views: (A) Data Table, (B) Main View, (C) Chat Box, and (D) Pattern View.
+As shown in Fig.
+7, the interface includes four main views: (A) Data Table, (B) Main View, (C) Chat Box, and (D) Pattern View.
 When a user uploads a table, the back-end of VisTR processes the tabular data into visualization references and stores them in the vector database, while the front-end interface displays the uploaded table in Data Table, and presents the visualization for the data in the Main View.
@@
 Additionally, Pattern View provides multimodal information (description text, different chart types, and candidate sketches) for a single variable’s typical change patterns, including descriptive text, various charts, and potential sketch candidates.
-6 Evaluation
+VI Evaluation
 To thoroughly evaluate VisTR, we conduct both quantitative evaluations and case studies.
 In the quantitative evaluations (Sect.
-6.1), we assess the alignment ability of our fine-tuned MLLM.
-Subsequently, we present two cases on actual application scenarios (Sects.
-6.2 & 6.3), representing a possible pathway through multimodal interactions to complete a time-series table reasoning task.
-6.1 Effectiveness of Multimodal Visualization Alignment
+VI-A), we assess the alignment ability of our fine-tuned MLLM.
+Subsequently, we present two classic cases on actual application scenarios (Sects.
+VI-B & VI-C), representing a possible pathway through multimodal interactions to complete a time-series table reasoning task.
+VI-A Effectiveness of Multimodal Visualization Alignment
 We evaluate the alignment performance of our multimodal visualization alignment model from two perspectives: quantitative evaluation on text-chart retrieval, and user recognition on sketch-chart retrieval.
-6.1.1 Baselines and Metrics
+VI-A1 Baselines and Metrics
 We compare our fine-tuned MLLM against several pre-trained models: CLIP[21], OpenCLIP[71], and UniChart[36].
@@
 WF adopts the weighted average considering the unevenness of the categories, which is calculated as: $WF=\frac{1}{T_{N}}\sum_{i=1}^{T_{N}}\frac{2\times P_{i}\times R_{i}}{(P_{i}+R_{i})}$, where $R_{i}$ and $P_{i}$ are the recall and precision of the $i^{th}$ trend-word category, respectively.
-6.1.2 Text-Chart Retrieval
+VI-A2 Text-Chart Retrieval
 To rigorously assess the performance of our fine-tuned visualization alignment model in aligning chart-text pairs, we ensure that the evaluation dataset includes chart types of bar, line, and area chart, as well as all $T_{N}=23$ categories of trend words.
 Specifically, we randomly select 20% of the data for each chart type and each trend category from the chart-text pairing dataset for evaluation.
-Table 1: Experimental results on chart-to-text retrieval test for all and individual chart types.
+TABLE I: Experimental results on chart-to-text retrieval test for all and individual chart types.
 Model All Bar Line Area
@@
 Results.
-The quantitative results are presented in Table 1.
+The quantitative results are presented in Table.
+I.
 Our fine-tuned MLLM demonstrates robust performance in aligning chart and text modalities, with overall accuracy and weighted F1-score exceeding 85% across all chart types.
@@
 Figure 8: Ratings of the visual similarity between pairs of the matched sketches and charts across our fine-tuned MLLM and three pre-trained models based on a 5-point Likert scale.
-6.1.3 Sketch-Chart Retrieval
+VI-A3 Sketch-Chart Retrieval
 We validate the alignment for chart-sketch pairs through a user study.
@@
 The ratings are gathered using a 5-point Likert scale[75], which provides a quantitative measure of the user agreement or disagreement with the similarities of chart-sketch pairs.
-As shown in Figure 7, our model achieved the highest similarity agreement among participants (Mean = 4.33, SD = 0.67), and the Unichart achieved the lowest agreement with Mean = 2.25 and SD = 1.17.
+As shown in Fig.
+8, our model achieved the highest similarity agreement among participants (Mean = 4.33, SD = 0.67), and the Unichart achieved the lowest agreement with Mean = 2.25 and SD = 1.17.
 Given the non-normal distribution of collected ratings, we initiated our analysis by conducting the Friedman test to evaluate the statistical significance of ratings among the four models.
@@
 To further pinpoint the specific differences among models, we employed the Wilcoxon signed-rank test for paired comparisons, accompanied by Bonferroni correction to control errors.
-The pairwise comparisons showed that our model significantly outperformed the others.
-The detailed comparison results are shown in Figure 7.
+The pairwise comparisons showed that our model significantly outperformed the others, as detailed in Fig.
+8.
 The poor performance of UniChart may be due to its focus on OCR features rather than chart features, leading to a weaker grasp of the varied patterns present within charts and hand-drawn sketches.
 The comparison results underscore the importance of encoding patterns in our alignment model, confirm the judicious choice of CLIP-ViT as our base pre-trained model, and highlight the superiority of our fine-tuned model in binding sketch and chart modalities.
-6.2 Case Study 1: Financial Data Exploration
+VI-B Case Study 1: Financial Data Exploration
 In this case, we consider a financial data exploration scenario.
@@
 The average retrieval time for querying a visualization reference stored in Chroma is within 0.6s.
-As shown in Figure 8, Mike explores the data as follows.
+As shown in Fig. 9, Mike explores the data as follows.
   • Table Summarization.
@@
 Figure 10: Case study 2: multivariate analysis for the uploaded air pollutant dataset. With VisTR, Emily recognizes typical data change patterns and relationships between different air pollutants.
-6.3 Case Study 2: Air Pollutant Analysis
+VI-C Case Study 2: Air Pollutant Analysis
 In the second case, we demonstrate the workflow of mining relationships between data change patterns of different variables.
@@
 In the backend, VisTR generates a total of 102,787 visualization references, of which 40,243 are stored in Chroma after pruning.
-Figure 9 illustrates Emily’s analysis process using VisTR to explore the multivariate data.
+Fig. 10 illustrates Emily’s analysis process using VisTR to explore the multivariate data.
   • Free Exploration.
@@
 In this manner, Emily gains a comprehensive insight into the multivariate air pollutant data, recognizing both the typical data change patterns and correlations among various air pollutant indicators.
-7 Discussion
+VII Discussion
 The case studies have demonstrated VisTR’s capability in recognizing and explaining data change patterns within tabular data.
@@
 Supporting a broader range of table types is feasible, leveraging methods such as subset embedding[33] and learning-based approaches[34].
-  • General user queries.
+  • General queries.
 VisTR focuses on tasks related to reasoning about data change patterns in tables.
@@
 To achieve a comprehensive exploration and reasoning of a data table, we require more general user queries, which may necessitate the incorporation of external knowledge sources.
+  • Dateset bias.
+The successful alignment of our model heavily relies on acquiring high-quality training data.
+However, the dataset used for fine-tuning, while demonstrating its effectiveness, is limited in size and diversity, posting challenges to generalizability.
+This limitation becomes particularly evident when handling complex or unconventional sketch representations, where a richer and more diverse dataset is essential for robust model fine-tuning.
+The dataset collection process itself introduces potential biases and inconsistencies.
+Scaling up the dataset to address these limitations presents significant challenges.
+First, the limited number of users contributing to user-labeled sketches may embed subjective preferences into the chart-sketch-text pairings.
+While crowd-sourcing strategies offer a way to accelerate data collection, they are prone to similar subjective biases and the quality control for such large-scale manual efforts remains a significant challenge.
+An innovative alternative lies in data augmentation techniques to reduce reliance on manual sketch collection.
+Second, the filtered trend words from textual descriptions may oversimplify visual information, neglecting nuances in semantics.
+To enrich textual descriptions, a potential solution proposed by Time-LLM[76] is to generate text prototypes by assembling candidate vocabularies.
+Then the generated rich text descriptions can be paired with charts through crowd-sourced rating frameworks[77].
   • Intent-enhanced sketches.
@@
 As a reason, VisTR offers insights into common patterns within the uploaded table.
-However, expressing complex and less common patterns solely through hand-drawn sketches can be challenging for users[32].
+However, expressing complex or uncommon patterns solely through hand-drawn sketches can be challenging for users[22].
 To address this challenge, several solutions can be considered.
-First, incorporating additional interactive features such as dragging, resizing, and erasing, sketches can provide users with greater flexibility in adjusting their intent[36].
+First, incorporating additional interactive features like dragging, resizing, and erasing can provide users with greater flexibility in adjusting their intent[78].
 Summarizing multi-attribute temporal patterns and introducing a more effective time-series segmentation method have shown promise in addressing this issue[79].
 By implementing these methods, we can further enhance the capabilities of VisTR to support sketches with enhanced intents.
-Scaling up the dataset.
-The successful alignment of our model heavily relies on acquiring high-quality training data.
-Scaling up the dataset to enhance the generalizability of VisTR presents significant challenges for the data augmentation and user labeling strategies currently employed.
-Collecting user-labeled sketches is time-consuming, and incorporating intents into sketches would further increase complexity.
-To address this limitation, one solution is to leverage crowdsourcing strategies, although this still requires significant manual effort.
-Another innovative technique would be to use large models to simulate and augment a variety of hand-drawn patterns.
-LLM-powered data augmentation has been a promising direction[11], and we are interested in using this approach for scaling up the dataset.
+VisTR vs.
+Traditional Methods.
+VisTR introduces an innovative attempt to convert time-series tables into visual representations, enabling efficient multi-modal binding and continuous context reasoning.
+Compared to traditional subsequence or sketch matching methods, such as metric-based[41], regular expression-based[74] or deep-learning-based approaches[72], VisTR processes varying data facet lengths into uniform vision tokens, offering smoother multi-modal interaction and significant resource savings due to reduced prompt lengths[80].
+However, visual representations come with notable limitations.
+They often underperform traditional metric-based methods in matching precision, incur higher storage costs, and may ignore shorter time subsequences in visualization referencing module.
+Additionally, fine-tuning models for aligning modalities of chart, sketch, and text is resource-intensive.
+Despite these challenges, the rapid advancements of multi-modal large language models (MLLMs) make this research direction promising.
+Future efforts should focus on improving precision and optimizing storage efficiency to fully leverage the potential of visual representations in advancing multi-modal frameworks.
 Potential scenarios.
@@
 Moreover, VisTR could also find applications in fraud detection, as visualizations generated from transactional data could help recognize suspicious patterns indicative of fraudulent activities.
-8 Conclusion
-This work introduces VisTR system, built upon a novel framework for table reasoning, expanding upon existing LLM-based table reasoning research.
-By leveraging visualizations as representations, the framework offers several advantages, including the ability to recognize data change patterns, mitigate LLM drift, and support visual-based exploration.
-The implementation of VisTR requires an effective multimodal visualization alignment model that binds chart, text, and sketch modalities together in a joint embedding space.
-This is accomplished by fine-tuning an MLLM with a new dataset of chart-text and chart-sketch pairings.
-Combined with strategies employed for visualization referencing, indexing, and pruning, we have developed an interactive system that facilitates multimodal interactions with time-series tables.
-The feasibility of VisTR is demonstrated through quantitative evaluations of the visualization alignment model and practical case studies on time-series data tables.
-This proposed framework offers a new research direction for table reasoning and enhances the fuzzy exploration process of existing LLM-based table reasoning methods.
-We aim to extend its support to a wider range of table types, accommodate more generalized user queries, and explore additional application scenarios in the future.
-Acknowledgements.
+VIII Conclusion
+We present VisTR, a new framework for time-series table reasoning that leverages visualizations as representations to bridge raw data and human cognition.
+By enabling enhanced pattern recognition, mitigating context drift, and supporting visual-based exploration, VisTR addresses key limitations of existing LLM-based methods.
+VisTR integrates a fine-tuned MLLM, trained on a newly constructed dataset of chart-text and chart-sketch pairs, to align charts, text, and sketches in a unified embedding space.
+Combined with strategies for visualization referencing, pruning, and indexing, VisTR provides an interactive system for seamless multimodal interactions with time-series tables.
+Quantitative evaluations and case studies demonstrate its effectiveness in aligning multimodal inputs and improving reasoning accuracy.
+VisTR opens new research directions for integrating visualizations into table reasoning and enhancing multimodal data exploration.
+Future work will extend its applicability to broader table types, generalized queries, and diverse application scenarios.